### v1

# Fair Contracts

[…]
ylred
Fair Contracts
Matteo Castiglioni
matteo.castiglioni@polimi.it
Politecnico di Milano
   Junjie Chen
junjchen9-c@my.cityu.edu.hk
Kyushu University
   Yingkai Li
yk.li@nus.edu.sg
National University of Singapore
We introduce and study the problem of designing optimal contracts under fairness constraints on the task assignments and compensations. We adopt the notion of envy-free (EF) and its relaxations, $\epsilon$-EF and envy-free up to one item (EF1), in contract design settings. Unlike fair allocations, EF contracts are guaranteed to exist. However, computing any constant-factor approximation to the optimal EF contract is NP-hard in general, even using $\epsilon$-EF contracts. For this reason, we consider settings in which the number of agents or tasks is constant. Notably, while even with three agents, finding an EF contract better than $\frac{2}{5}$ approximation of the optimal is NP-hard, we are able to design an FPTAS when the number of agents is constant, under relaxed notions of $\epsilon$-EF and EF1. Moreover, we present a polynomial-time algorithm for computing the optimal EF contract when the number of tasks is constant. Finally, we analyze the price of fairness in contract design. We show that the price of fairness for exact EF contracts can be unbounded, even with a single task and two agents. In contrast, for EF1 contracts, the price of fairness is bounded between $\Omega(\sqrt{n})$ and $O(n^{2})$, where $n$ is the number of agents.
1 Introduction
Contract theory is one of the fundamental fields in microeconomics. It tackles the challenge of moral hazard due to agents’ unobservable costly actions (Grossman and Hart,, 1992; Bolton and Dewatripont,, 2004). Recently, it has attracted attention from the computer science community since Dütting et al., (2019). Recent developments have primarily focused on the computational challenges in finding the optimal contracts (e.g., Guruganesh et al.,, 2021; Castiglioni et al.,, 2025; Dütting et al.,, 2023), the approximation guarantees (e.g., Dütting et al.,, 2019; Alon et al.,, 2023; Ezra et al.,, 2024) or learnabilities (e.g., Bacchiocchi et al.,, 2023; Chen et al.,, 2024). See the survey by Dütting et al., (2024) for a more comprehensive discussion of the recent developments.
In this paper, we investigate fairness concerns in contract design. Motivated by recent findings from the World Bank highlighting stalled progress in reducing the global prosperity gap (The World Bank Group,, 2024), we consider how a principal can allocate tasks and payments to agents in a way that is both profit-maximizing and fair. As inequality continues to rise, ensuring equitable treatment in labor markets—especially in algorithmically mediated platforms such as Uber, Didi, or Amazon Mechanical Turk—has become an increasingly important challenge (Chipriyanov et al.,, 2024). Moreover, fair treatment is often mandated by regulatory frameworks (e.g., Employment and Social Affairs Committee,, 2019), requiring platforms to ensure nondiscrimination and transparency in task assignment and compensation. We model this challenge through the lens of algorithmic contract design, where a principal assigns tasks and compensations to agents to maximize revenue while satisfying fairness constraints. Our goal is to understand the structural and computational implications of imposing fairness, and to develop algorithms that achieve (approximately) optimal profit subject to fairness guarantees.
1.1 Model
We study the design of fair contracts in a setting with $m$ tasks and $n$ agents. All tasks are costly, and each agent needs to pay some cost for accomplishing any given task. In this setting, each task is assigned to exactly one agent, while each agent may receive multiple tasks. For each task $j\in[m]$, the principal also provides a linear contract $\alpha_{j}\in[0,1]$ to incentivize the agent to accomplish the task. A main constraint in our contract design setting is that the task allocations and the linear contracts are fair. In particular, we adopt the well-known solution concept, envy freeness (EF), as the fairness constraint. In the context of fair contracts, envy freeness implies that agents prefer their own assigned sets of tasks and associated linear contracts over others. The objective of the designer is to design the task allocations and contracts to maximize the profit subject to fairness constraints.
A trivial fair contract is to provide an arbitrary task allocation and a linear contract $\alpha_{j}=0$ for all task $j$. This is equivalent to not assigning any task to any agent. To circumvent such a trivial case, we employ the no free disposal assumption like the literature in fair division (e.g., Amanatidis et al.,, 2023), where all task $j$ must be assigned to some agent, and the linear contract $\alpha_{j}\in[0,1]$ must be sufficient to incentivize the agent to accomplish the task. We term the contract satisfying this assumption as the full-allocation contract. Under the mild assumption that for all task $j$, there exists an agent such that he can be incentivized to accomplish task $j$ with a linear contract $\alpha_{j}\in[0,1]$, which obviously is also necessary for the existence of the full-allocation contract, we show that an envy-free full-allocation contract always exists. This is a sharp contrast to the classic fair division literature where envy-free allocation may not exist.
Although envy-free contracts always exist, computing a profit-maximizing envy-free contract remains computationally challenging. This difficulty arises from the inherent tension between fairness and revenue extraction. For example, consider a setting with a single task and two agents: one generates low revenue but accepts a low payment, while the other generates significantly higher revenue but requires a slightly higher compensation. The revenue-maximizing choice would be to assign the task to the latter. However, this may induce envy from the first agent, who would prefer the higher-paying task. To ensure envy-freeness, the principal must instead assign the task to the lower-value agent, sacrificing profit. In this paper, we formalize this trade-off and study the algorithmic complexity of computing (approximately) optimal contracts under fairness constraints. We provide both hardness results and efficient algorithms in various settings.
1.2 Our Results
Table 1: Summary of computational results. All results are derived with respect to the benchmark of optimal EF contracts, except for those stated explicitly otherwise.
Contracts General settings Settings with constant number of agents Settings with constant number of tasks
EF contracts Inapproximable within $\frac{1}{c}$ for any $c\geq 1$ Inapproximable within factor $\frac{2}{5}$ for $3$ agents Polynomially solvable
NP-hard for $2$ agents
$\epsilon$-EF contracts Inapproximable within $\frac{1}{c}$ for any $c\geq 1$, even for a constant $\epsilon>0$ Additive FPTAS Polynomially solvable w.r.t. optimal $\epsilon$-EF contracts
EF1 contracts Inapproximable within $\frac{7}{10}$ w.r.t. optimal EF1 contracts Inapproximable within $\frac{7}{10}$ for $3$ agents w.r.t. optimal EF1 contracts Polynomially solvable w.r.t. optimal EF1 contracts
Additive FPTAS
A key contribution of this paper is the introduction of a natural and tractable contract design model that captures the interplay between fairness constraints and profit maximization. As an illustration, we analyze the problem of computing (approximately) optimal envy-free (EF) contracts and their relaxations. Our main algorithmic and hardness results are summarized in Table 1. In addition to algorithmic considerations, we study the price of fairness, quantifying the loss in profit incurred by imposing fairness constraints. Below, we present a detailed overview of our results.
We first show that in general settings with an arbitrary number of agents and tasks, there is a strong negative result that it is impossible to approximate the EF optimal contracts within any constant factor in polynomial time. This is true even if we relax the fairness constraint to $\epsilon$-EF (Definition 2.6 and Theorem 3.2). The high-level idea of the hardness result is to construct a reduction from an independent set problem where the size of the independent set is either very small or very large (Alon et al.,, 1995; Trevisan,, 2001).
The hardness results in general settings motivate us to investigate the computational questions in more restrictive settings with a constant number of agents or a constant number of tasks. For the relatively easy case with a constant number of tasks, we first observe that the computation of the optimal envy-free contract can be solved in polynomial time via linear programming if the task assignment is exogenously given. With a constant number of tasks, the number of possible assignments is polynomial in the input size. Therefore, finding the optimal envy-free contract is polynomial-time solvable in this case (Proposition 4.1). This result also immediately implies efficient algorithms for finding optimal $\epsilon$-EF or EF1 contracts.
The case with a constant number of agents is more challenging. We show that even with only $3$ agents, it is NP-hard to approximate the optimal EF contracts better than some constant factor (Theorem 3.4). This rules out the possibility of a multiplicative PTAS for EF contracts. The proof of the hardness result is built upon a reduction from the partition problem (Garey and Johnson,, 2002). Interestingly, the hardness results can be easily extended to approximating the optimal EF1 contracts (Theorem B.1). Moreover, the problem remains NP-hard for settings even with two agents (Proposition 3.5).
We circumvent the challenges by considering the computation of $\epsilon$-EF or EF1 contracts with respect to the benchmark of optimal EF contracts, and adopt the objective of additive approximation instead of multiplicative approximation. We provide an additive FPTAS for both $\epsilon$-EF and EF1 contracts through carefully designed dynamic programming (Theorems 4.3 and 4.4). The dynamic programming for EF1 contracts is non-trivial, which requires a novel adaptive grid for discretization in order to ensure that the EF1 constraints can always be satisfied without any error. Interestingly, the satisfaction of EF1 constraints is indeed retained by transforming the almost EF constraints (i.e., $\epsilon$-EF) with an additive specially-designed error to almost EF constraints (i.e., EF1) in a multiplicative sense.
Finally, we investigate the price of fairness in contract design settings, i.e., the impact of introducing fairness constraints on the principal’s profit. For exact EF constraints, the price of fairness could be arbitrarily high, meaning that the principal’s profit could be arbitrarily worse than that without fairness constraints. This could happen even in the cases with two agents. We then examine the price of fairness for two relaxations of fairness constraints. Specifically, we show that with EF1 constraints, the price of fairness is bounded between $\Omega(\sqrt{n})$ and $O(n^{2})$. Another relaxation we consider is motivated by the recent fair division literature on subsidies (e.g., Elmalem et al.,, 2025). We propose envy-free with subsidies (EFS) contracts, where the principal can compensate each agent with additional state-independent transfers. We show that although EFS contracts could reduce the price of fairness in some cases, the worst-case price of fairness remains unbounded.
1.3 Related Work
In this section, we provide discussions on related works. In particular, our paper is closely related to contract design and fair division.
Related works on algorithmic contracts.
Contract theory is a foundational area in microeconomics, with a rich literature dating back several decades (Holmstrom,, 1980; Grossman and Hart,, 1992; Bolton and Dewatripont,, 2004; Carroll,, 2015; Castro-Pires et al.,, 2024). More recently, contract design has attracted growing interest from the computer science community, where algorithmic methods enable the analysis of richer and more complex settings. The algorithmic study of contracts was initiated by Babaioff et al., (2006). Subsequent work explored both the complexity of optimal contracts (Dutting et al.,, 2021) and approximation guarantees for simple linear contracts (Dütting et al.,, 2019). A line of recent work integrates moral hazard with adverse selection, providing algorithmic solutions in both single-parameter (Alon et al.,, 2021, 2023) and multi-parameter settings (Guruganesh et al.,, 2021; Castiglioni et al.,, 2021, 2022; Guruganesh et al.,, 2023; Castiglioni et al.,, 2025). Beyond this, other recent directions include agent-designed contracts (Bernasconi et al.,, 2024) and contract design jointly with information structure (Castiglioni and Chen,, 2025). We refer readers to the recent survey by Dütting et al., (2024) for a comprehensive overview of algorithmic contract design.
Our contract design setting is closely related to the literature on combinatorial contracts(e.g., Dütting et al.,, 2023, 2022; Dutting et al.,, 2024; Deo-Campo Vuong et al.,, 2024; Ezra et al.,, 2024; Duetting et al.,, 2025). This line of work primarily focuses on single-task settings, addressing the combinatorial challenges arising from multiple agents or multiple actions. A notable exception is the recent work of Alon et al., (2025), who consider multi-task environments and provide a polynomial-time algorithm for computing a constant-factor approximation to the optimal contract. A key difference between our model and that of Alon et al., (2025) lies in the structure of task allocation. In their setting, each agent may exert effort for at most one task. In contrast, we consider the setting where each task must be assigned to exactly one agent, and agents may exert effort across multiple tasks. This introduces a different combinatorial structure and leads to new algorithmic challenges.
Despite extensive research in algorithmic contract theory, contract design under fairness constraints has received limited attention. Fehr et al., (2007) conducted experimental studies showing that, when fairness concerns are present, bonus contracts can outperform standard incentive contracts. More recently, Dasaratha et al., (2023); Tłuczek et al., (2025) consider contract design with equity constraints. Their model is orthogonal to ours: they consider a single-task setting and restrict the compensation scheme to be identical across agents. In contrast, we study a multi-task environment, where the principal simultaneously assigns tasks and designs individualized incentive contracts to maximize profit, subject to envy-free constraints.
Related works on fair division.
Our work relates to the literature on fair division under envy-free constraints. Envy freeness is a fairness notion that has been studied extensively in the literature (e.g., Beynier et al.,, 2019; Foley,, 1966; Lipton et al.,, 2004). There are several well-known relaxations of envy freeness, including envy-free up to one item (EF1) (Budish,, 2011), envy-free up to any item (EFX) (Caragiannis et al.,, 2019), weighted-envy-freeness (WEF) (Chakraborty et al.,, 2021) and proportionality (Steinhaus,, 1948; Suksompong,, 2016).
Our work also conceptually relates to the fair division of chores (e.g., Heydrich and van Stee,, 2015; Dehghani et al.,, 2018). In the fair division of chores, each agent has some cost (or disutility in the literature) for a chore, and all agents are forced to complete the assigned chores. However, in our model, rational agents may choose to shirk from the chores, and the principal needs to provide additional monetary incentives for them to complete the chores.
Finally, the price of fairness in our paper relates to the literature on the price of fairness in good and chore allocation (Bei et al.,, 2021; Bu et al.,, 2025). For example, Bei et al., (2021) proves a lower bound of $\Omega(\sqrt{n})$ and an upper bound of $O(n)$ for the price of EF1. Our price of fairness with subsidies relates to the fair division with subsidies (Halpern and Shah,, 2019; Elmalem et al.,, 2025). For more details, we refer the interested readers to the survey by Liu et al., (2024).
2 Model
Setting.
A principal has a set of tasks $\mathcal{M}$ to complete and can delegate them to a set of agents $\mathcal{N}$. Let $m=|\mathcal{M}|$ and $n=|\mathcal{N}|$ be some finite values. In particular, the principal has to assign each task $j\in\mathcal{M}$ to an agent $i\in\mathcal{N}$. If task $j\in\mathcal{M}$ is assigned to agent $i$, then agent $i\in\mathcal{N}$ can choose to exert effort $c_{i,j}\in[0,1]$ to try to complete the task, which then succeeds with probability $p_{i,j}\in[0,1]$. If the task $j$ is completed, the principal gets reward $r_{j}\in[0,1]$; otherwise, the reward is zero. If the agent $i$ chooses not to exert effort, task $j$ will never succeed.
To incentivize the agents to work, the principal designs a contract for each task. In this paper, we focus on the simple but widely adopted class of linear contracts. Specifically, a linear contract determines the fraction $\alpha_{j}\in[0,1]$ of the reward given to the agent to which the task is assigned, if the task is completed. Under the linear contract $\alpha_{j}$, if an agent $i$ successfully completes task $j$, it will receive payment $\alpha_{j}\cdot r_{j}$; otherwise, it will not receive any payment.
Fair Contracts.
The main question we examine in this paper is to design allocations of tasks to agents which satisfy fairness constraints. We focus on the simple case in which each task can be assigned to at most one agent, while each agent may receive the delegation of multiple tasks. We use $S=\{S_{i}\}_{i\in\mathcal{N}}$ to denote the allocation, where $S_{i}\subseteq\mathcal{M}$ is the set of tasks assigned to agent $i$. Then, allocation $S$ is feasible if $S_{i}\cap S_{j}=\emptyset$ for all $i\neq j$. A contract $(S,\alpha)$ in our setting consists of a feasible allocation $S$ and the linear contracts $\alpha_{j}$ for each task $j$.
As a starting point for the problem of fair contracts, we consider the fairness notion of envy-freeness, which has been extensively studied in the fair division literature. With that, we define the envy-free (EF) contracts.
Definition 2.1 (Envy-free Contracts).
A contract $(S,\alpha)$ is envy-free if
$\sum_{k\in S_{i}}\max\{\alpha_{k}p_{i,k}r_{k}-c_{i,k},0\}\geq\sum_{k\in S_{j}}\max\{\alpha_{k}p_{i,k}r_{k}-c_{i,k},0\},\quad\forall j\neq i.$ (1)
Intuitively, given the contract $(S,\alpha)$, envy-free contracts require that each agent $i$ prefers their assigned set of tasks over switching to another agent’s set of tasks and exerting optimal effort.
To align with the literature on fair division, where all items must be assigned to one of the agents, we consider a setting in which the principal is restricted to full allocation contracts. That is, every task must be assigned to some agent, ensuring that no tasks are left unallocated. Moreover, we require that agents always exert effort to try to complete the tasks. [^1] Notice that without this further restriction, assigning a task to an agent that does not exert effort is essentially equivalent to not assigning the task at all.
Definition 2.2 (Full Allocation Contracts).
A contract $(S,\alpha)$ is a full allocation contract if $\cup_{i\in\mathcal{N}}S_{i}=\mathcal{M}$ and
$\alpha_{j}p_{i,j}r_{j}-c_{i,j}\geq 0,\quad\forall i\in\mathcal{N},j\in S_{i}.$ (IR constraints)
Throughout our paper, we maintain the restriction to full allocation contracts. Under this restriction, the envy-free constraints (1) can be simplified as follows:
$\sum_{k\in S_{i}}\alpha_{k}p_{i,k}r_{k}-c_{i,k}\geq\sum_{k\in S_{j}}\max\{\alpha_{k}p_{i,k}r_{k}-c_{i,k},0\},\quad\forall j\neq i.$ (EF constraints)
We make the following mild assumption to ensure the existence of full allocation contracts.
Assumption 2.3.
For any task $j\in\mathcal{M}$, there exists an agent $i\in\mathcal{N}$ such that $p_{i,j}r_{j}-c_{i,j}\geq 0$.
Intuitively, if the social welfare of a task is negative for every possible agent that performs the task, not assigning this task is the correct approach. Hence, we can simply ignore the task in the allocation process. Moreover, Assumption 2.3 is reasonable in practice. For example, on a ride-hailing platform, if a customer sending a request is located in a too rural region or the payment is too low, the platform will have no incentives to consider that request and instead put the customer into a long waiting queue.
Existence of EF Contracts.
Before moving forward, we first show two preliminary observations about the EF full allocation contracts. Indeed, under 2.3, it is without loss of optimality to focus on envy-free full allocation contracts.
Proposition 2.4.
Under 2.3, for any envy-free contract $(S,\alpha)$, there exists an envy-free full allocation contract $(\hat{S},\hat{\alpha})$ that generates a weakly higher revenue for the principal.
Proof.
Suppose that in an envy-free contract $(S,\alpha)$, task $j$ is unallocated. Among the set of agents $\mathcal{M}_{j}=\{i\mid p_{i,j}r_{j}-c_{i,j}\geq 0\}$, assign task $j$ to the agent $i^{*}\in\mathcal{M}_{j}$ where $i^{*}=\operatorname*{arg\,min}_{i}\frac{c_{i,j}}{p_{i,j}r_{j}}$ and set the contract $\alpha_{j}=\frac{c_{i^{*},j}}{p_{i^{*},j}r_{j}}$. Same operations are applied to other unallocated tasks. Obviously, these modifications do not change the IR and EF constraints but increase the principal’s revenue. ∎
Therefore, from now on, we will refer to the full allocation contracts simply as contracts. More surprisingly, in sharp contrast to the fair allocation literature, where envy-free allocations may not exist, we show that envy-free (full allocation) contracts always exist.
Proposition 2.5.
There always exists a (full allocation) contract that is envy free, and there exists an algorithm that finds it in polynomial time.
Proof.
One possible EF contracts is constructed as follows: for every task $j$, among the set of agents $\mathcal{M}_{j}=\{i\mid p_{i,j}r_{j}-c_{i,j}\geq 0\}$, assign task $j$ to the agent $i^{*}\in\mathcal{M}_{j}$ where $i^{*}=\operatorname*{arg\,min}_{i}\frac{c_{i,j}}{p_{i,j}r_{j}}$ and set the contract $\alpha_{j}=\frac{c_{i^{*},j}}{p_{i^{*},j}r_{j}}$. This allocation is EF and the IR constraints are satisfied. ∎
The Principal’s Problem.
Given any contract $(S,\alpha)$, the expected revenue of the principal is
${\rm Rev}(S,\alpha)=\sum_{i\in\mathcal{N}}\sum_{k\in S_{i}}(1-\alpha_{k})p_{i,k}r_{k}.$
In this paper, the objective of the principal is to design a fair contract that maximizes the expected revenue, subject to the EF constraints and IR constraints. Specifically, we have the following optimization program.
$\max_{\{\alpha_{k}\}_{k\in\mathcal{M}},\{S_{i}\}_{i\in\mathcal{N}}}\quad$ ${\rm Rev}(S,\alpha)$ (OPT-EF)
subject to (IR constraints), (EF constraints)
$S_{i}\cap S_{j}=\emptyset,\quad\forall i\neq j$
$\bigcup_{i\in\mathcal{N}}S_{i}=\mathcal{M}.$
We also use $\rm{OPT\text{-}EF}$ to denote the optimal objective value for program (OPT-EF).
Relaxed Notions of EF contracts.
We also consider two relaxations of EF contracts: $\epsilon$-EF contracts and EF1 contracts. $\epsilon$-EF contracts relax the EF constraints by some $\epsilon>0$, while EF1 contracts relax EF to envy-free up to one task. Note that Proposition 2.4 and 2.5 also hold for $\epsilon$-EF contracts and EF1 contracts.
Definition 2.6 ($\epsilon$-Envy-Free Contracts).
A contract $(S,\alpha)$ is $\epsilon$-envy-free if
$\sum_{k\in S_{i}}\alpha_{k}p_{i,k}r_{k}-c_{i,k}\geq\sum_{k\in S_{j}}\max\{\alpha_{k}p_{i,k}r_{k}-c_{i,k},0\}-\epsilon,\quad\forall j\neq i.$ ($\epsilon$-EF constraints)
Definition 2.7 (Envy-Free-Up-to-One (EF1) Contracts).
A contract $(S,\alpha)$ is envy-free up to one task (EF1) if for agents $i,j$, there exists a task $k_{i,j}\in S_{j}$ such that
$\sum_{k\in S_{i}}\alpha_{k}p_{i,k}r_{k}-c_{i,k}\geq\sum_{k\in S_{j}\backslash\{k_{i,j}\}}\max\{\alpha_{k}p_{i,k}r_{k}-c_{i,k},0\},\quad\forall j\neq i.$ (EF1 constraints)
[…]
Additional notations.
Throughout the remaining sections of the paper, we let $\lceil x\rceil_{\mathcal{X}}$ be the smallest element in $\mathcal{X}$ greater than or equal to $x$, given a $x\in\mathbb{R}$ and a discrete set $\mathcal{X}$. Also, let $[n]\triangleq\{1,2,\dots,n\}$ for any integer $n$, and $\lfloor x\rfloor$ be the smallest integer greater than or equal to $x$.
3 Computational Hardness for EF Contracts
In this section, we show that despite the desirable property that envy-free contracts always exist (Proposition 2.5), finding a revenue-optimal one is computationally challenging. Specifically, we show that even finding an approximately optimal envy-free contract is NP-hard for any constant approximation factor (Theorem 3.2). Moreover, we show that even if we restrict our attention to settings with a constant number of agents, finding a contract with an approximation factor better than $\frac{5}{2}$ remains NP-hard (Theorem 3.4). The missing proofs from this section are provided in Section A.1.
3.1 Hardness for General Settings
In this section, we prove the hardness results of computing optimal EF contracts. In particular, we provide a stronger result that it is NP-hard to approximate optimal EF contracts within any constant factor, even using the relaxed $\epsilon$-EF contracts for approximation. The proof is by a reduction from the following problem.
[…]
  • there exists an independent set of size at least $\eta|V|$,
  • all the independent sets have size at most $\mu\eta|V|$,
determine which of the two cases holds.
[…]
Theorem 3.2.
For any constant $c\geq 1$, there exist constant $\epsilon>0$ such that, assuming P $\neq$ NP, there does not exist a polynomial-time algorithm that computes an $\epsilon$-EF contract with revenue at least a $\frac{1}{c}$ fraction of $\rm{OPT\text{-}EF}$.
Proof.
[…]
In the following, we let $\gamma=|V|$, $\beta=|E|$, and we define the constant $\delta=\frac{1}{8ck^{2}}$, and $\epsilon=\delta/3$. Notice that since the maximum degree is $k$, we have $\beta\leq\frac{\gamma k}{2}$.
To prove the statement, we show that if there exists an independent set of size at least $\eta|V|$, then there exists an EF contract with principal’s utility at least $\frac{1}{2}\eta\gamma$. Moreover, if all independent set have size at most $\mu\eta\gamma$, then all $\epsilon$-EF contracts achieve principal utility at most $\frac{1}{2c}\eta\gamma$. This is enough to prove the statement.
Construction.
Given a graph $G=(V,E)$, we construct an instance as follows. We first define the set of agents. For each edge $e\in E$, there exists an agent $b_{e}$. Additionally, we have one “efficient” agent $b^{*}$. Then, we define two sets of tasks:
  • For each vertex $v\in V$, we have a task $d_{v}$ with reward $r_{d_{v}}=1$. The probabilities of success and costs are constructed as follows.
    • Let $p_{b^{*},d_{v}}=1$ and $c_{b^{*},d_{v}}=1/2$.
    • If node $v\in e$, i.e., the edge $e$ connects $v$ to another node, then let $p_{b_{e},d_{v}}=\delta$ and $c_{b_{e},d_{v}}=0$.
    • Let $p_{b,d_{v}}=0$ and $c_{b,d_{v}}=0$ for all other agents $b$.
  • For each edge $e\in E$, we define a task $d_{e}$ with $r_{d_{e}}=\delta/2$. The probabilities of success and costs are defined as follows.
    • For agent $b_{e}$, let $p_{b_{e},d_{e}}=1$ and $c_{b_{e},d_{e}}=0$.
    • Let $p_{b,d_{e}}=c_{b,d_{e}}=0$ for all other agents $b$.
Intuitively, the principal would like to assign as many tasks $d_{v}$, $v\in V$, as possible to agent $b^{*}$. However, suppose that we assign two tasks $d_{v}$ and $d_{v^{\prime}}$ to $b^{*}$, where $(v,v^{\prime})=e\in E$. Notice that the linear contract on each of this task is at least $1/2$ by IR. Hence, agent $b_{e}$ would get $\delta$ performing tasks $d_{v}$ and $d_{v}^{\prime}$. Thus, it must be the case that the payment to agent $b_{e}$ from $d_{e}$ (the only other task with positive success probability) is $\delta$, which is impossible since $r_{d_{e}}=\delta/2$.
Sufficiency.
Assume that there is an independent set $I$ such that $|I|\geq\eta\gamma$.
Consider the set of tasks $\{d_{v}\}_{v\in V}$. We assign all the tasks $d_{v}$ with $v\in I$ to agent $b^{*}$ and we set contract $\alpha_{d_{v}}=\frac{1}{2}$ for every $v\in I$. Apparently, the IR constraints for agent $b^{*}$ are satisfied. We assign every remaining task $d_{v}$, where $v\in V\setminus I$, to any agent $b_{e^{*}}$ with $e^{*}\in E$. We set the contract for these tasks $d_{v}$ to $\alpha_{d_{v}}=0$. Note that this will not increase the utility of agent $b_{e^{*}}$, but it will increase the principal’s revenue by at most $(\gamma-|I|)\delta$.
Next, we consider the set of tasks $\{d_{e}\}_{e\in E}$. For any edge $e\in E$, note that at most one of its nodes $v\in e$ belongs to the set $I$. If such a node $v\in e\cap I$ exists, then agent $b_{e}$ would receive utility $\frac{1}{2}\delta$ by switching to the set of tasks $I$ assigned to $b^{*}$. Hence, we assign the task $d_{e}$ to this agent $b_{e}$ by setting contract $\alpha_{d_{e}}=1$, which leads to satisfying EF constraints. Otherwise, i.e., if $e\cap I=\emptyset$, the corresponding agent $b_{e}$ gets utility 00 by switching to agent $b^{*}$’s assigned set $I$. Hence, we can assign the task $d_{e}$ to agent $b^{*}$ by setting contract $\alpha_{d_{e}}=0$. Note that this will not change agents’ utilities, and the EF constraints continue to hold. The principal gets revenue 00 for this set of tasks.
Hence, we conclude that the principal’s revenue is
$\frac{1}{2}|I|+(\gamma-|I|)\delta\geq\frac{1}{2}\eta\gamma.\qed$
Necessity.
Assume that all the independent sets have size at most $\gamma\eta\mu$.
Consider an allocation $S$ and contracts $\mu$. Let $S_{b^{*}}$ be the set of tasks assigned to $b^{*}$ and $V^{*}$ be the subset of nodes in $V$ whose corresponding tasks are assigned to agent $b^{*}$, i.e., $V^{*}=\{v\in V\mid d_{v}\in S_{b^{*}}\}$. By the IR constraints, it must be the case that $\alpha_{d_{v}}\geq\frac{1}{2}$ for all $v\in V^{*}$.
We first show that the size $|V^{*}|\leq\gamma\eta\mu$. The following argument is by contradiction. Suppose $|V^{*}|>\gamma\eta\mu$. By assumption, we have that $V^{*}$ is not an independent set. Therefore, there must exist some edge $e=(v_{1},v_{2})$ such that both nodes $v_{1},v_{2}\in V^{*}$. This implies that both tasks $d_{v_{1}}$ and $d_{v_{2}}$ are assigned to agent $b^{*}$. Hence, by the EF constraints, if an agent $b_{e}$ switches to work on the set of tasks $S_{b^{*}}$, it would receive payments at least
$\alpha_{d_{v_{1}}}p_{b_{e},d_{v_{1}}}r_{d_{v_{1}}}+\alpha_{d_{v_{2}}}p_{b_{e},d_{v_{2}}}r_{d_{v_{2}}}\geq\delta.$ (2)
Hence, the payment to agent $b_{e}$ is at least $\delta$. However, since we are restricted to linear contracts, the payments to agent $b_{e}$ is upperbounded by $\delta/2$. Indeed, the only other task in which agent $b_{e}$ is $d_{e}$ with payment at most $\delta/2$. Hence, the EF constraint is violated by at least $\delta/2>\epsilon$ by Equation (2), reaching a contradiction.
Thus, we can restrict to allocations in which $|V^{*}|\leq\gamma\eta\mu$. In this case, the principal’s revenue is at most
$\frac{1}{2}|V^{*}|+\delta(\gamma-|V^{*}|)+\frac{\delta}{2}\beta\leq\frac{1}{2}\eta\mu\gamma+\delta(\gamma+\frac{\gamma k}{2})\leq\frac{1}{2c}\eta\gamma$
where in the first inequality we use $\beta\leq\frac{\gamma k}{2}$. This concludes the proof.
This immediately implies that there is no PTAS for computing EF-optimal contracts.
Corollary 3.3.
For any constant $c\geq 1$, assuming P $\neq$ NP, there does not exist a polynomial-time algorithm that computes an EF contract with revenue at least a $\frac{1}{c}$ fraction of $\rm{OPT\text{-}EF}$.
3.2 Hardness for Settings with A Constant Number of Agents
In this section, we show that even in the restricted setting with a constant number of agents, computing an EF contract with an approximation ratio better than $\frac{5}{2}$ is still NP-hard. This rules out the possibility of a multiplicative PTAS for EF contracts. In Section B.1, we extend the analysis to show that computing an EF1 with an approximation ratio better than $\frac{10}{7}$ is also NP-hard. Moreover, in Section B.2, we show that for any $\epsilon<\frac{1}{5}$, computing an $\epsilon$-EF with an approximation ratio better than $\frac{20\epsilon+6}{10}$ is NP-hard.
Theorem 3.4.
When there are three agents, assuming P $\neq$ NP, there does not exist a polynomial-time algorithm that computes an envy-free contract with revenue at least a $\frac{2}{5}$ fraction of $\rm{OPT\text{-}EF}$.
Proof.
Our proof is by a reduction from the partition problem: given a set of $\ell$ integers $n_{1},n_{2},\dots,n_{\ell}$, find a subset of integers $\mathcal{S}$ such that $\sum_{j\in\mathcal{S}}n_{j}=\sum_{j\in[\ell]}n_{j}/2$.
Construction.
We construct an instance where there are $3$ agents, i.e., $n=3$. Given a set of integers $n_{1},n_{2},\dots,n_{\ell}$, we define the instance as follows. There exists a task $j^{*}$ whose reward is $r_{j^{*}}=1$. The agents’ success probabilities for task $j^{*}$ are $p_{1,j^{*}}=1$, $p_{2,j^{*}}=p_{3,j^{*}}=\frac{1}{10}$, and their costs are $c_{1,j^{*}}=\frac{1}{2}$, $c_{2,j^{*}}=c_{3,j^{*}}=0$.
Let $M=10\sum_{j\in[\ell]}n_{j}$. For each $n_{j}$, we define a task $j$ with reward $r_{j}=n_{j}/M$. Additionally, we define the agents’ success probabilities as $p_{1,j}=0$, $p_{2,j}=p_{3,j}=1$ and costs as $c_{1,j}=c_{2,j}=c_{3,j}=0$.
Sufficiency.
We show that if there exists a partition, then the principal’s revenue is at least $\frac{1}{2}$.
Suppose that there is a set of indexes $\mathcal{S}$ such that $\sum_{j\in\mathcal{S}}n_{j}=\sum_{j\in[\ell]}n_{j}/2$. Consider the following contract. Assign task $j^{*}$ to agent $1$ with contracts as $\alpha_{j^{*}}=\frac{1}{2}$. Moreover, assign all the tasks in $\mathcal{S}$ to agent $2$, and all the other tasks to agent $3$. We set the contracts $\alpha_{j}=1$ to all the tasks $j\in[\ell]$.
We are left to show that the allocation satisfies the IR constraint and it guarantees revenue at least $\frac{1}{2}$. First, the allocation is envy-free. Indeed, by the construction above, agent $1$ receives utility 00. Due to $p_{1,j}=0$ for all $j\in[\ell]$, agent $1$ receives utility 00 switching to the tasks assigned to one of the other agents. Agent $2$ receives utility
$\sum_{j\in\mathcal{S}}r_{j}p_{2,j}-c_{2,j}=\sum_{j\in\mathcal{S}}\frac{n_{j}}{M}=\frac{1}{20}.$
Switching to the tasks assigned to agent $1$, agent $2$ gets utility $\alpha_{j^{*}}r_{j^{*}}p_{2,j^{*}}=\frac{1}{20}$. In addition, switching to the task assigned to agent $3$, agent $2$ gets utility $\sum_{j\in[\ell]\setminus\mathcal{S}}\alpha_{j}r_{j}p_{2,j}-c_{2,j}=\frac{1}{20}$ since $\mathcal{S}$ is a partition. Hence, the agent $2$ does not envy other agents. A similar argument applies to agent $3$.
We conclude noticing that the principal gets positive revenue only from task $j^{*}$ and in particular $\frac{1}{2}$.
Necessity.
We prove that if the partition does not exists, the principal’s revenue is at most $\frac{1}{5}$.
We prove the result by contradiction. Assume that the principal’s revenue is strictly greater than $\frac{1}{5}$. Notice that if agent $1$ is not assigned the task $j^{*}$, the principal’s revenue is at most:
$r_{j^{*}}p_{2,j^{*}}+\sum_{j\in[\ell]}r_{j}p_{2,j}=\frac{1}{10}+\sum_{j\in[\ell]}\frac{n_{j}}{M}=\frac{1}{5},$
which is obtained, for instance, by assigning all the tasks to agent $2$.
Hence, task $j^{*}$ must be assigned to agent $1$ since by assumption the principal’s revenue is strictly greater than $\frac{1}{5}$. This directly implies that the payment for task $j^{*}$ is at least $\alpha_{j^{*}}\geq\frac{1}{2}$ by the IR constraint of agent $1$.
Now, we analyze the EF constraints of agent $2$ and $3$. Since agent $2$ does not envy agent $1$, the set of tasks $S_{2}\subseteq[\ell]$ assigned to agent $2$ must guarantee utility:
$\sum_{j\in S_{2}}\alpha_{j}r_{j}p_{2,j}=\sum_{j\in S_{2}}\alpha_{j}\frac{n_{j}}{M}\geq\alpha_{j^{*}}r_{j^{*}}p_{2,j^{*}}\geq 1/20.$
where the first inequality is by the EF constraints with respect to agent $1$, and the second inequality is by $\alpha_{j^{*}}\geq\frac{1}{2}$. Therefore, we have that
$\sum_{j\in S_{2}}\frac{n_{j}}{10\sum_{j\in[\ell]}n_{j}}=\sum_{j\in S_{2}}\frac{n_{j}}{M}\geq\sum_{j\in S_{2}}\alpha_{j}\frac{n_{j}}{M}\geq\frac{1}{20}$
which further implies that
$\sum_{j\in S_{2}}n_{j}\geq\frac{1}{2}\sum_{j\in[\ell]}n_{j}.$
By similar arguments, we have that for the set of tasks $S_{3}\subseteq[\ell]$ assigned to agent $3$ guarantees:
$\sum_{j\in S_{3}}n_{j}\geq\frac{1}{2}\sum_{j\in[\ell]}n_{j}.$
Hence, it must be the case that $\sum_{j\in S_{2}}n_{j}=\sum_{j\in S_{3}}n_{j}=\frac{1}{2}\sum_{j\in[\ell]}n_{j}$, which implies that $S_{2}$ defines a partition, leading to a contradiction.
we conclude the proof noticing that the gap is $\frac{1}{5}/\frac{1}{2}=\frac{2}{5}$ as promised. ∎
Moreover, in the more extreme case with only two agents, we show that computing the optimal EF contract remains NP-hard. It remains an open problem if good approximations can be obtained for the two-agent case.
Proposition 3.5.
When there are only two agents, assuming P $\neq$ NP, there does not exist a polynomial-time algorithm that computes the optimal envy-free contract.
4 Polynomial-time Algorithms in Restricted Settings
In this section, we focus on the design of polynomial-time algorithms that may achieve “good” revenue in restricted settings. We first show that in the case with a constant number of tasks, the optimization program can be directly formulated as a set of linear programs with polynomial size, and hence the optimal solution can be found in polynomial time (Proposition 4.1).
Recall the strong negative results Theorem 3.2 that in general settings of an arbitrary number of agents, there is no PTAS for approximating the optimal EF contracts even with $\epsilon$-EF contracts. This motivates the second part of our results, studying settings with a constant number of agents under relaxed notions of EF. Though Theorem 3.4 shows that the problem doesn’t admit a PTAS under exact EF unless P=NP, such hardness results can be circumvented by considering two relaxations, $\epsilon$-EF (Definition 2.6) and EF1 (Definition 2.7). For each relaxation, we provide an (additive) FPTAS via dynamic programming (Theorems 4.3 and 4.4).
4.1 Polynomial-time Algorithms for a Constant Number of Tasks
We first observe from Program (OPT-EF) that if the allocation of tasks is fixed, Program (OPT-EF) can be solved through linear programming. Therefore, if the number of tasks $m$ is a constant, we can enumerate all the possible allocations of tasks, where the number of such allocations is at most $O(n^{m})$. For each allocation, we solve the relative linear program. All these steps clearly take polynomial time. Hence, we immediately have the following result.
Proposition 4.1.
When the number of tasks is a constant, there exists a polynomial-time algorithm that computes an optimal envy-free contract in ${\rm poly}(n,m)$ time.
The proof of Proposition 4.1 is in Appendix A.2. Note that this immediately implies poly-time algorithms for optimal $\epsilon$-EF and EF1 contracts (in Appendix B.3).
4.2 FPTAS for a Constant Number of Agents
In light of the hardness result for exact envy-free contracts (Theorem 3.4), we focus on polynomial-time algorithms for finding approximately optimal contracts under relaxed notions of fairness. We start by introducing a general dynamic programming template in Section 4.2.1. Then, we design an FPTAS for $\epsilon$-EF contracts in Section 4.2.2. Finally, we show how to generalize the idea to design an FPTAS for EF1 contracts in Section 4.2.3.
4.2.1 A general dynamic programming template
In this section, we show that once we discretize the set of possible contracts and utilities, optimal contracts can be computed efficiently. In particular, we consider a setting in which
  • the contract $\hat{\alpha}_{j}$ belong to a discrete set $\mathcal{D}_{j}\subset[0,1]$ for each task $j\in\mathcal{M}$.
  • the utility $\hat{U}_{i,j}(\hat{\alpha}_{j})$ of an agent $i$ performing a task $j$ under contract $\hat{\alpha}_{j}\in\mathcal{D}_{j}$ belongs to a discrete set $\mathcal{U}_{i}\subset[0,1]$.
  • the principal’s utility $\hat{U}^{P}(\hat{\alpha}_{j},i)$ from a task $j$ performed by agent $i$ under contract $\hat{\alpha}_{j}$ belongs to a discrete set $\mathcal{H}\subset[0,1]$.
To design an FPTAS for $\epsilon$-EF contracts, we will use the standard uniform grid. However, the generality of the framework will be essential to find EF1 allocations.
The pseudocode of our algorithm is provided in Algorithm 1. The algorithm cycles over all the tasks $\ell\in\mathcal{M}$ and keeps track of the possible utility vectors that can be induced by an IR contract. Formally, it keeps track of a set $\mathcal{L}_{\ell}$ of all the possible inducible tuples $(\hat{S},\hat{\alpha},\hat{h},(\hat{v}_{i,j})_{i,j\in\mathcal{N}}))$, where
  • $\hat{S}$ is a partial allocation of the first $\ell$ tasks;
  • $\hat{\alpha}$ includes the contracts assigned to the first $\ell$ tasks;
  • $\hat{h}$ represents the principal’s utility from the first $\ell$ tasks;
  • $\hat{v}_{i,j}$ is the utility of an agent $i$ working on the tasks assigned to agent $j$ belonging to the first $\ell$ tasks.
The inducible tuples are computed as follows. For each contract $\hat{\alpha}_{\ell}\in\mathcal{D}_{\ell}$ and agent $k\in\mathcal{N}$, it verifies if such contract is IR for agent $k$ on task $\ell$. If it is the case, it computes all the partial allocations obtained by assigning task $\ell$ to agent $k$ with contract $\hat{\alpha}_{\ell}$ in all the possible partial allocations $(\hat{S},\hat{\alpha},\hat{h},(\hat{v}_{i,j})_{i,j\in\mathcal{N}})\in\mathcal{L}_{\ell-1}$ (see Lines from 1 to 1), where we compute the discrete utilities of the principal and agents after allocating task $\ell$ as follows:
$h^{\prime}=\hat{h}+\hat{U}^{P}(\hat{\alpha}_{\ell},k)$ (3a)
$v^{\prime}_{i,k}=\hat{v}_{i,k}+\hat{U}_{i,\ell}(\hat{\alpha}_{\ell})\quad\forall i\in\mathcal{N}$ (3b)
$v^{\prime}_{i,j}=\hat{v}_{i,j}\quad\forall i\in\mathcal{N},\forall j\neq k$ (3c)
For simplicity, let $y=(\hat{h},(\hat{v}_{i,j})_{i,j\in\mathcal{N}})$ be the utility profile of agents and the principal. Note that our goal is only to have a representative contract (and allocation) for each inducible utility profile $y$. If a couple $(\hat{S}^{\prime},\hat{\alpha}^{\prime})$ induces the same utility profile as another couple, it is removed (See Line 1).
Finally, for each utility profile $y$, the algorithm returns a couple $(\hat{S},\hat{\alpha})$ that induces it. Formally,
Proposition 4.2.
Algorithm 1 returns a set $\mathcal{C}$ such that:
  • all $(\hat{S},\hat{\alpha})\in\mathcal{C}$ are IR
  • For each $(S,\alpha)$ with $\alpha\in\times_{i\in\mathcal{M}}\mathcal{D}_{i}$ that is IR, there exists an $(\hat{S},\hat{\alpha})\in\mathcal{C}$ such that
    $\sum_{i\in\mathcal{N}}\sum_{j\in S_{i}}\hat{U}^{P}(\alpha_{j},i)=\sum_{i\in\mathcal{N}}\sum_{j\in\hat{S}_{i}}\hat{U}^{P}(\hat{\alpha}_{j},i)$
    $\sum_{k\in S_{j}}\hat{U}_{i,k}(\alpha_{k})=\sum_{k\in\hat{S}_{j}}\hat{U}_{i,k}(\hat{\alpha}_{k})\quad\forall i,j\in\mathcal{N}.$
Proof.
All contracts in $\mathcal{C}$ are IR thanks to Line 1. Then, the remaining of the proof follows from standard dynamic programming arguments. ∎
5 $\mathcal{L}_{0}\leftarrow\{(\emptyset^{n},\emptyset,0^{nm+1})\}$;
6 for $\ell=1,\dots,m$ do
7       $\mathcal{L}_{\ell}\leftarrow\emptyset$ ;
8       for $\hat{\alpha}_{\ell}\in\mathcal{D}_{\ell}$ do
9             for agent $i=1,\dots,n$ do
10                   if $\hat{\alpha}_{\ell}p_{i,\ell}r_{\ell}-c_{i,\ell}<0$  then
11                         continue;
12
13                   end if
14                   for  $(\hat{S},\hat{\alpha},y)\in\mathcal{L}_{\ell-1}$ do
15                         Let $\hat{S}_{i}^{\prime}=\hat{S}_{i}\cup\{\ell\}$ ;
16                         Let $\hat{S}_{j}^{\prime}=\hat{S}_{j}$ for each $j\neq i$ ;
17                         Let $\hat{\alpha}^{\prime}=(\hat{\alpha},\hat{\alpha}_{\ell})$;
18                         Compute $y^{\prime}$ as per Equation (3);
19                         if  it does not exists $(S,\alpha,y)\in\mathcal{L}_{\ell}$ where $y=y^{\prime}$  then
20                               $L_{\ell}=L_{\ell}\cup\{(\hat{S}^{\prime},\hat{\alpha}^{\prime},y^{\prime})\}$ ;
21
22                         end if
23
24                   end for
25
26             end for
27
28       end for
29
30 end for
return $\mathcal{C}=\{\hat{S},\hat{\alpha}:\textnormal{there exists a $y$ with}(\hat{S},\hat{\alpha},y)\in\mathcal{L}_{m}\}$.
2 Input: $\{r_{j}\}_{j\in\mathcal{M}}$, $\{p_{i,j}\}_{i\in\mathcal{N},j\in\mathcal{M}}$, $\{c_{i,j}\}_{i\in\mathcal{N},j\in\mathcal{M}}$, $m,n$, discretization of contracts $(\mathcal{D}_{j})_{j\in\mathcal{M}}$, of agent’s utility $(\mathcal{U}_{i})_{i\in\mathcal{N}}$, and principal’s utility $\mathcal{H}$, utility function $\hat{U}_{i,j}(\cdot)$ and $\hat{U}^{P}(\cdot,\cdot)$
Output: set of couples (allocation,contract) $\mathcal{C}$.
Algorithm 1 Dynamic programming to find all utilities profiles
4 Input: $\{r_{j}\}_{j\in\mathcal{M}}$, $\{p_{i,j}\}_{i\in\mathcal{N},j\in\mathcal{M}}$, $\{c_{i,j}\}_{i\in\mathcal{N},j\in\mathcal{M}}$, $m,n$, discretization of contracts $(\mathcal{D}_{j})_{j\in\mathcal{M}}$, of agent’s utility $(\mathcal{U}_{i})_{i\in\mathcal{N}}$, and principal’s utility $\mathcal{H}$, utility function $\hat{U}_{i,j}(\cdot)$ and $\hat{U}^{P}(\cdot,\cdot)$
Output: set of couples (allocation,contract) $\mathcal{C}$.
The running time of the algorithm depends on the discretizations, which will be analyzed later when we instantiate it for $\epsilon$-EF contracts or EF1 contracts in the following sections. In particular, it depends on the possible utility tuples that can be induced.
4.2.2 $\epsilon$-EF Contracts
Our first result gives an FPTAS algorithm for $\epsilon$-EF contracts.
Theorem 4.3.
When the number of agents is a constant, there is an algorithm that computes an $\epsilon$-envy-free contract with expected revenue at least $\rm{OPT\text{-}EF}-\epsilon$ in time polynomial in the instance size and $\frac{1}{\epsilon}$.
Proof.
Let denote an optimal solution to Program (OPT-EF) with $\alpha^{*}=(\alpha_{1}^{*},\alpha_{2}^{*},\dots,\alpha_{m}^{*})$, and its corresponding fair allocations of tasks $S^{*}=(S_{1}^{*},S_{2}^{*},\dots,S_{n}^{*})$.
The proof proceeds as follows. We build a discretized instance of the problem, we solve it with dynamic programming, and we show that one of the solutions returned by Algorithm 1 is approximately EF and approximately optimal.
Discretized instance.
We start building the discretized instance. Given a parameter $\epsilon>0$, we define $\delta=\frac{\epsilon}{m}$. We discretize the set of contract and principal/agent utilities with a grid of step $\delta$. Formally, we define
$\mathcal{D}_{i}=\{0,\delta,\ldots,1\}\quad\forall i\in\mathcal{M}$ (4)
$\mathcal{U}_{i}=\{0,\delta,\ldots,1\}\quad\forall i\in\mathcal{N}$
=$\mathcal{H}=\{0,\delta,\ldots,1\}$
In the following since all sets $\mathcal{D}_{i}$ and $\mathcal{U}_{i}$ are equal for $i\in\mathcal{N}$, we will refer to these sets simply with $\mathcal{D}$ and $\mathcal{U}$, respectively. Finally, we define
$\hat{U}_{i,j}(\hat{\alpha}_{j})=\lceil\hat{\alpha}_{j}p_{i,j}r_{j}-c_{i,j}\rceil_{\mathcal{U}}\quad\forall i\in\mathcal{N},j\in\mathcal{M},\hat{\alpha}_{j}\in\mathcal{D}$ (5)
$\hat{U}^{P}(\hat{\alpha}_{j},i)=\lceil(1-\hat{\alpha}_{j})p_{i,j}r_{j}\rceil_{\mathcal{H}}\quad\forall i\in\mathcal{N},j\in\mathcal{M},\hat{\alpha}_{j}\in\mathcal{D}$
Existence of a good solution with discretized contracts
Consider the contract $\bar{\alpha}$ obtained setting $\bar{\alpha}_{i}=\lceil\alpha^{*}_{i}\rceil_{\mathcal{D}}$. We show that $(S^{*},\bar{\alpha})$ is almost optimal and almost EF. Formally:
$\begin{split}\sum_{k\in S^{*}_{i}}\bar{\alpha}_{k}p_{i,k}r_{k}-c_{i,k}&\geq\sum_{k\in S^{*}_{i}}\alpha_{k}^{*}p_{i,k}r_{k}-c_{i,k}\\ &\geq\sum_{k\in S^{*}_{j}}\max\{\alpha_{k}^{*}p_{i,k}r_{k}-c_{i,k},0\}\\ &\geq\sum_{k\in S^{*}_{j}}\max\{(\bar{\alpha}_{k}-\delta)p_{i,k}r_{k}-c_{i,k},0\}\\ &\geq\sum_{k\in S^{*}_{j}}\max\{\bar{\alpha}_{k}p_{i,k}r_{k}-c_{i,k},0\}-\epsilon\quad\quad\forall i,j\in\mathcal{N}\end{split}$ (6)
where the second inequality is due to $(\alpha^{*},S^{*})$ satisfying the EF constraints. Moreover, the IR constraints for the tasks continue to hold as we round up the contracts.
Finally, we get that the principal’s revenue under the new (approximate) contracts $(\bar{\alpha},S^{*})$ is at least $\rm{OPT\text{-}EF}-\epsilon$, which is
$\sum_{i\in\mathcal{N}}\sum_{k\in S^{*}_{i}}(1-\bar{\alpha}_{k})p_{i,k}r_{k}\geq\sum_{i\in\mathcal{N}}\sum_{k\in S^{*}_{i}}(1-{\alpha}_{k}^{*}-\delta)p_{i,k}r_{k}\geq\sum_{i\in\mathcal{N}}\sum_{k\in S^{*}_{i}}(1-{\alpha}_{k}^{*})p_{i,k}r_{k}-\epsilon=\rm{OPT\text{-}EF}-\epsilon.$ (7)
Find a good solution with discretize contracts.
The pseudocode of our algorithm is in Algorithm 2. Our algorithm enumerates over all the possible couples $(S,\alpha)\in\mathcal{C}$ returned by Algorithm 1, and outputs the $3\epsilon$-EF contract which maximizes the principal’s utility. Let denote this contract with $(\tilde{S},\tilde{\alpha})$.
By Proposition 4.2, Algorithm 1 find a set $\mathcal{C}$ which includes an allocation $\hat{S}$ and a contract $\hat{\alpha}$ that induce the same discretized utility profile as $(\bar{\alpha},S^{*})$. Formally, we have that
$\sum_{i\in\mathcal{N}}\sum_{j\in S_{i}^{*}}\hat{U}^{P}(\bar{\alpha}_{j},i)=\sum_{i\in\mathcal{N}}\sum_{j\in\hat{S}_{i}}\hat{U}^{P}(\hat{\alpha}_{j},i)$
$\sum_{k\in S_{j}^{*}}\hat{U}_{i,k}(\bar{\alpha}_{k})=\sum_{k\in\hat{S}_{j}}\hat{U}_{i,k}(\hat{\alpha}_{k})\forall i,j\in\mathcal{N}$
$\hat{\alpha}_{j}p_{i,j}r_{j}-c_{i,j}\geq 0\quad\forall i\in\mathcal{N},j\in\hat{S}_{i}.$
The last step of the proof shows that $(\tilde{S},\tilde{\alpha})$ achieves principal’s utility at least $\rm{OPT\text{-}EF}-2\epsilon$. Indeed, this contract is $3\epsilon$-EF by construction, while all the contracts in $\mathcal{C}$ are IR. To do so, it is sufficient to show that $(\hat{S},\hat{\alpha})$ is $3\epsilon$-EF and $2\epsilon$-optimal, since this contract belongs to $\mathcal{C}$.
We start by showing that the contract is approximately EF. Take two agents $i$ and $j$. Then
$\sum_{k\in\hat{S}_{i}}\hat{\alpha}_{k}p_{i,k}r_{k}-c_{i,k}$ $\geq\sum_{k\in\hat{S}_{i}}(\hat{U}_{i,k}(\hat{\alpha}_{k})-\delta)$
$\geq\sum_{k\in\hat{S}_{i}}\hat{U}_{i,k}(\hat{\alpha}_{k})-\epsilon$
$=\sum_{k\in S^{*}_{i}}\hat{U}_{i,k}(\bar{\alpha}_{k})-\epsilon$
$\geq\sum_{k\in S^{*}_{i}}\bar{\alpha}_{k}p_{i,k}r_{k}-c_{i,k}-\epsilon$
$\geq\sum_{k\in S^{*}_{j}}\max\{\bar{\alpha}_{k}p_{i,k}r_{k}-c_{i,k},0\}-2\epsilon$
$\geq\sum_{k\in S^{*}_{j}}\hat{U}_{i,k}(\bar{\alpha})-3\epsilon$
$=\sum_{k\in\hat{S}_{j}}\hat{U}_{i,k}(\hat{\alpha})-3\epsilon$
$\geq\sum_{k\in\hat{S}_{j}}\hat{\alpha}_{k}p_{i,k}r_{k}-c_{i,k}-3\epsilon.$
where the second and fourth inequalities are by the definition of $\hat{U}_{i,k}$, the fifth inequality is by (6) and the equalities are by Proposition 4.2. Moreover, it is $2\epsilon$-optimal since
$\sum_{i\in\mathcal{N}}\sum_{k\in\hat{S}_{i}}(1-\hat{\alpha}_{k})p_{i,k}r_{k}$ $\geq\sum_{i\in\mathcal{N}}\sum_{k\in\hat{S}_{i}}U^{P}(\hat{\alpha}_{k},i)-\epsilon$
$=\sum_{i\in\mathcal{N}}\sum_{k\in S^{*}_{i}}U^{P}(\bar{\alpha}_{k},i)-\epsilon$
$\geq\sum_{i\in\mathcal{N}}\sum_{k\in S^{*}_{i}}(1-\bar{\alpha}_{k})p_{i,k}r_{k}-\epsilon$
$\geq\rm{OPT\text{-}EF}-2\epsilon.$
where the last inequality is by (7).
We conclude the proof noticing that the algorithm runs in polynomial time. Indeed, Algorithm 2 clearly run in polynomial time excluding the calls to Algorithm 1. Moreover, it easy to see that each call to Algorithm 1 runs in polynomial time since the possible vectors of agents/principal’s cumulative utilities have polynomial size (each utility belongs to $\{0,\delta,\ldots,m\}$. ∎
5 $\delta\leftarrow\frac{\epsilon}{m}$ ;
6 Define $\mathcal{D}_{i}$, $\mathcal{U}_{i}$, $\mathcal{H}$ as per Equation (4) and functions $\hat{U}_{i,j}(\cdot)$ and $\hat{U}^{P}(\cdot,\cdot)$ (Eq. (5);
7 $\mathcal{C}\leftarrow$ Run Algorithm 1 with discretization $\mathcal{D}_{i}$, $\mathcal{U}_{i}$, $\mathcal{V}_{i}$, $\mathcal{H}$, $\mathcal{V}^{P}$ and functions $\hat{U}_{i,j}(\cdot)$, $\hat{U}^{P}(\cdot,\cdot)$;
8 $\mathcal{C}^{*}\leftarrow\emptyset$ ;
9 for  $(S,\alpha)\in\mathcal{C}$ do
10       if $(S,\alpha)$ satisfies $3\epsilon$-EF  then
11             $\mathcal{C}^{*}=\mathcal{C}^{*}\cup\{{S},\alpha\}$ ;
12
13       end if
14
15 end for
return $({\tilde{\alpha}},{\tilde{S}})\in\mathcal{C}^{*}$ with maximum principal utility.
2 Input: $\epsilon>0$ , $\{r_{j}\}_{j\in\mathcal{M}}$, $\{p_{i,j}\}_{i\in\mathcal{N},j\in\mathcal{M}}$, $\{c_{i,j}\}_{i\in\mathcal{N},j\in\mathcal{M}}$, $m,n$
Output: The contracts $(\tilde{\alpha},\tilde{S})$.
Algorithm 2 Computing $3\epsilon$-EF Contracts
4 Input: $\epsilon>0$ , $\{r_{j}\}_{j\in\mathcal{M}}$, $\{p_{i,j}\}_{i\in\mathcal{N},j\in\mathcal{M}}$, $\{c_{i,j}\}_{i\in\mathcal{N},j\in\mathcal{M}}$, $m,n$
Output: The contracts $(\tilde{\alpha},\tilde{S})$.
4.2.3 EF1 Contracts
The analyses for EF1 contract is much more involved. In particular, approximating the agents’ utility with an additive grid incurs an additive error in the EF constraint, which does not directly provide any EF1 guarantees.
Our approach revolves around getting allocations which provide a multiplicative approximation of the EF constraints, and the relation between multiplicative approximation of EF and EF1. Formally, we prove the following.
Theorem 4.4.
When the number of agents is a constant, there is an algorithm that computes an EF1 contract with expected revenue at least $\rm{OPT\text{-}EF}-\epsilon$ in time polynomial in the instance size and $\frac{1}{\epsilon}$.
Proof.
We will use the template of the $\epsilon$-EF dynamic programming approach designed in the previous section, enhanced with an initial guessing step and an adaptive grid. The pseudocode of our algorithm is in Algorithm 3.
Denote the optimal solution of Program (OPT-EF) as $\alpha^{*}=(\alpha_{1}^{*},\alpha_{2}^{*},\dots,\alpha_{m}^{*})$, and its corresponding fair allocations of tasks as $S^{*}=(S_{1}^{*},S_{2}^{*},\dots,S_{n}^{*})$. Given the desired approximation $\epsilon>0$ for the principal’s revenue, we set errors $\nu=\min\{\epsilon,\frac{1}{6m}\}$ and $\delta=\frac{\nu}{m}$.
Characterize the range of utilities.
Since, given an optimal allocation $S^{*}$, the optimal contract is a solution to an LP, standard arguments imply that the utility of each agent is 00 or at least $2^{-f(I)}$, where $f(I)$ is a polynomial function of the instance size (Bertsimas and Tsitsiklis,, 1997).
Guessing step.
We consider the following set of approximate agent’s utilities:
$\mathcal{G}=\{0\}\cup\{m2^{-i}\}_{i\in\{0,1,\ldots,f(I)+\log(m)\}}.$
Then, we enumerate over all the tuples $(U_{i})_{i\in\mathcal{N}}\in\mathcal{G}^{n}$ which include a guess of the utility for each agent. Notice that there exists a guess $(U_{i})_{i\in\mathcal{N}}$ such that:
$\sum_{k\in S^{*}_{i}}\alpha_{k}^{*}p_{i,k}r_{k}-c_{i,k}\in[U_{i}/2,U_{i}]\quad\forall i\in\mathcal{N}.$ (8)
In the following, we focus on the execution of Algorithm 3 when the guess is correct and hence we assume that Equation (8) holds.
Adaptive grid.
Once we fix the utility profile $(U_{i})_{i\in\mathcal{N}}$, we can define an adaptive grid for the dynamic programming as a function of the guesses.
We define a set of possible discrete contracts as follows. Consider a task $j\in\mathcal{M}$ and an agent $i\in\mathcal{N}$. Let $\underline{\alpha}_{i,j}$ be the maximum $\alpha$ in the set
$\{\alpha\in[0,1]:\alpha p_{i,j}r_{j}-c_{i,j}\leq U_{i}\}.$ (9)
and let $\underline{\alpha}_{j}=\min_{i}\underline{\alpha}_{i,j}$. Then, we define the set
$\mathcal{D}_{i,j}=\begin{cases}\{c_{i,j}/(r_{j}p_{i,j})+\delta i(\underline{\alpha}_{j}-c_{i,j}/(r_{j}p_{i,j}))\}_{i\in\{0,1,\ldots,1/\delta\}}\quad\quad\textnormal{if }\underline{\alpha}_{j}\geq c_{i,j}/(r_{j}p_{i,j})\\ \mathcal{D}_{i,j}=\emptyset\hskip 182.09746pt\textnormal{otherwise}\end{cases}$ (10)
Notice that for each $j\in\mathcal{M}$, at least one of the sets $\{\mathcal{D}_{i,j}\}_{i\in\mathcal{N}}$ is not empty thanks to Assumption 2.3. Finally, we set the possible contracts for a task $j\in\mathcal{M}$ as
$\mathcal{D}_{j}=\cup_{i\in\mathcal{N}}\mathcal{D}_{i,j}.$ (11)
Now, we define the set of possible discretized utilities with a normalized uniform grid. In particular, for each agent $i$, we define the discretized utilities as
$\mathcal{U}_{i}=\{\delta kU_{i}\}_{k\in\{0,1,\ldots,1/\delta\}}\quad\forall i\in\mathcal{N}$ (12)
Finally, we use a uniform grid over the principal’s utility:
=$\mathcal{H}=\{0,\delta,\ldots,1\}$ (13)
Now, we are left to define the approximate utility functions $\hat{U}_{i,j}(\cdot)$, $\hat{U}^{P}(\cdot,\cdot)$. We let
$\hat{U}_{i,j}(\hat{\alpha}_{j})=\lceil\hat{\alpha}_{j}p_{i,j}r_{j}-c_{i,j}\rceil_{\mathcal{U}_{i}}\quad\forall i\in\mathcal{N},j\in\mathcal{M},\hat{\alpha}_{j}\in\mathcal{D}_{j}$ (14)
$\hat{U}^{P}(\hat{\alpha}_{j},i)=\lceil(1-\hat{\alpha}_{j})p_{i,j}r_{j}\rceil_{\mathcal{H}}\quad\forall i\in\mathcal{N},j\in\mathcal{M},\hat{\alpha}_{j}\in\mathcal{D}_{j}$
It is easy to see that $\lceil\hat{\alpha}_{j}p_{i,j}r_{j}-c_{i,j}\rceil_{\mathcal{U}_{i}}$ are well-defined thanks to the definition of $\mathcal{D}_{j}$. Moreover, it holds
$\hat{U}_{i,j}(\hat{\alpha}_{j})\in[\max\{\hat{\alpha}_{j}p_{i,j}r_{j}-c_{i,j},0\},\max\{\hat{\alpha}_{j}p_{i,j}r_{j}-c_{i,j},0\}+\delta U_{i}]\quad\forall i\in\mathcal{N},j\in\mathcal{M},\hat{\alpha}_{j}\in\mathcal{D}_{j}$ (15)
Existence of a good solution with discretized contracts.
We start showing that there exists a discretized contract that is almost EF and almost optimal. Consider the contract $\bar{\alpha}$ obtained by setting $\bar{\alpha}_{j}=\lceil\alpha^{*}_{j}\rceil_{\mathcal{D}_{j}}$ for each $j\in\mathcal{M}$. First of all, we have to show that such $\bar{\alpha}$ is well-defined. To do that, we have to guarantee that $\mathcal{D}_{j}$ includes at least one contract greater than $\alpha^{*}_{j}$ for each $j\in\mathcal{M}$.
Lemma 4.5.
$\bar{\alpha}$ is well-defined, i.e., $\alpha^{*}_{j}\leq\max_{\alpha_{j}\in\mathcal{D}_{j}}\alpha_{j}$ for all $j\in\mathcal{M}$.
Proof.
Consider a $j\in\mathcal{M}$. We show that $\alpha^{*}_{j}\leq\underline{\alpha}_{j}$. This is sufficient to show that $\bar{\alpha}_{j}$ is well-defined by the definition of $\mathcal{D}_{j}$.
We consider two cases: $\underline{\alpha}_{j}=1$ and $\underline{\alpha}_{j}<1$.
If $\underline{\alpha}_{j}=1$, it immediately implies $\alpha^{*}_{j}\leq\underline{\alpha}_{j}$. Hence, in the following arguments, we focus on the second case $\underline{\alpha}_{j}<1$.
Suppose by contradiction that $\alpha^{*}_{j}>\underline{\alpha}_{j}$. Let $i$ be such that $\underline{\alpha}_{j}=\underline{\alpha}_{i,j}$. Then, by definition,
$\underline{\alpha}_{j}=\frac{c_{i,j}+U_{i}}{r_{j}p_{i,j}},$ (16)
where we use $\underline{\alpha}_{j}<1$. Furthermore, suppose that $i^{\prime}$ is the agent such that $j\in S^{*}_{i^{\prime}}$. Then, it holds:
$U_{i}$ $\geq\sum_{k\in S^{*}_{i}}\alpha^{*}_{k}p_{i,k}r_{k}-c_{i,k}$
$\geq\sum_{k\in S^{*}_{i^{\prime}}}\max\{\alpha^{*}_{k}p_{i,k}r_{k}-c_{i,k},0\}$
$\geq\alpha_{j}^{*}p_{i,j}r_{j}-c_{i,j}$
$>\underline{\alpha}_{j}p_{i,j}r_{j}-c_{i,j}$
$=\frac{U_{i}+c_{i,j}}{p_{i,j}r_{j}}p_{i,j}r_{j}-c_{i,j}$
$\geq U_{i},$
where the first inequality is by Equation (8), the second inequality is by EF and the equality follows from Equation (16). Hence, we reach a contradiction. ∎
Now, we show that $(S^{*},\bar{\alpha})$ is almost optimal and almost EF. Formally:
Lemma 4.6.
$(S^{*},\bar{\alpha})$ is IR and guarantees:
$\begin{split}\sum_{k\in S^{*}_{i}}\bar{\alpha}_{k}p_{i,k}r_{k}-c_{i,k}\geq\sum_{k\in S^{*}_{j}}\max\{\bar{\alpha}_{k}p_{i,k}r_{k}-c_{i,k},0\}-\nu U_{i}\quad\forall i,j\in\mathcal{N},\end{split}$ (17)
and
$\sum_{i\in\mathcal{N}}\sum_{k\in S^{*}_{i}}(1-\bar{\alpha}_{k})p_{i,k}r_{k}\geq\rm{OPT\text{-}EF}-\nu.$ (18)
Proof.
First, we prove that
$\max\{\bar{\alpha}_{j}p_{i,j}r_{j}-c_{i,j},0\}\leq\max\{{\alpha}^{*}_{j}p_{i,j}r_{j}-c_{i,j},0\}+\delta U_{i}\quad\forall i\in\mathcal{N},j\in\mathcal{M}.$ (19)
Consider an agent $i$ and a task $j$. We consider three cases:
  1. (i)
    $\underline{\alpha}_{j}\geq c_{i,j}/(r_{j}p_{i,j})$ and ${\alpha}^{*}_{j}\leq c_{i,j}/(r_{j}p_{i,j})$
  2. (ii)
    $\underline{\alpha}_{j}\geq c_{i}/(r_{j}p_{i,j})$ and ${\alpha}^{*}_{j}>c_{i}/(r_{j}p_{i,j})$
  3. (iii)
    $\underline{\alpha}_{j}<c_{i}/(r_{j}p_{i,j})$.
(i) If $\underline{\alpha}_{j}\geq c_{i,j}/(r_{j}p_{i,j})$ and ${\alpha}^{*}_{j}\leq c_{i,j}/(r_{j}p_{i,j})$, then both $\max\{\bar{\alpha}_{j}p_{i,j}r_{j}-c_{i,j},0\}$ and $\max\{{\alpha}^{*}_{j}p_{i,j}r_{j}-c_{i,j},0\}$ are 00, as $\bar{\alpha}_{j}=\lceil\alpha^{*}_{j}\rceil_{\mathcal{D}_{j}}\leq c_{i,j}/(r_{j}p_{i,j})$ by that the set $\mathcal{D}_{i,j}$ starts from $c_{i,j}/(r_{j}p_{i,j})$.
(ii) If $\underline{\alpha}_{j}\geq c_{i}/(r_{j}p_{i,j})$ and ${\alpha}^{*}_{j}\geq c_{i}/(r_{j}p_{i,j})$, then
$\max\{\bar{\alpha}_{j}p_{i,j}r_{j}-c_{i,j},0\}$ $=\bar{\alpha}_{j}p_{i,j}r_{j}-c_{i,j}$
$\leq({\alpha}^{*}_{j}+\delta(\underline{\alpha}_{j}-\frac{c_{i,j}}{p_{i,j}r_{j}}))p_{i,j}r_{j}-c_{i,j}$
$\leq{\alpha}^{*}_{j}p_{i,j}r_{j}+\delta(\underline{\alpha}_{j}p_{i,j}r_{j}-c_{i,j})-c_{i,j}$
$\leq{\alpha}^{*}_{j}p_{i,j}r_{j}-c_{i,j}+\delta U_{i}$
$=\max\{{\alpha}^{*}_{j}p_{i,j}r_{j}-c_{i,j},0\}+\delta U_{i},$
where the last inequality follows from the definition of $\underline{\alpha}_{i,j}$ in Equation (9) and $\underline{\alpha}_{j}\leq\underline{\alpha}_{i,j}$.
(iii) If $\underline{\alpha}_{j}<c_{i}/(r_{j}p_{i,j})$, then $\max\{\bar{\alpha}_{j}p_{i,j}r_{j}-c_{i,j},0\}$ and $\max\{{\alpha}^{*}_{j}p_{i,j}r_{j}-c_{i,j},0\}$ are 00, as $\bar{\alpha}_{j}=\lceil\alpha^{*}_{j}\rceil_{\mathcal{D}_{j}}\leq\underline{\alpha}_{j}$ by Lemma 4.5 and the definition of $\underline{\alpha}_{j}$. This concludes the proof of Equation (19).
Then, we use Equation (19) to prove Equation (17). Consider two agents $i$ and $j$. It holds:
$\sum_{k\in S^{*}_{i}}\bar{\alpha}_{k}p_{i,k}r_{k}-c_{i,k}$ $\geq\sum_{k\in S^{*}_{i}}\alpha_{k}^{*}p_{i,k}r_{k}-c_{i,k}$
$\geq\sum_{k\in S^{*}_{j}}\max\{\alpha_{k}^{*}p_{i,k}r_{k}-c_{i,k},0\}$
$\geq\sum_{k\in S^{*}_{j}}(\max\{\bar{\alpha}_{k}p_{i,k}r_{k}-c_{i,k},0\}-\delta U_{i})$
$\geq\sum_{k\in S^{*}_{j}}\max\{\bar{\alpha}_{k}p_{i,k}r_{k}-c_{i,k},0\}-\nu U_{i}\quad\forall i,j\in\mathcal{N},$
where the third inequality is by Equation (19). Moreover, the IR constraints for the tasks continue to hold as we round up the contracts.
Finally, we get that the principal’s revenue under the new approximate contracts $(\bar{\alpha},S^{*})$ is at least $\rm{OPT\text{-}EF}-\nu$, which is
$\sum_{i\in\mathcal{N}}\sum_{k\in S^{*}_{i}}(1-\bar{\alpha}_{k})p_{i,k}r_{k}\geq\sum_{i\in\mathcal{N}}\sum_{k\in S^{*}_{i}}(1-{\alpha}_{k}^{*}-\delta)p_{i,k}r_{k}\geq\sum_{i\in\mathcal{N}}\sum_{k\in S^{*}_{i}}(1-{\alpha}_{k}^{*})p_{i,k}r_{k}-\nu=\rm{OPT\text{-}EF}-\nu.$
Find a good solution with discretized contracts.
Our algorithm enumerates over all the possible utilities $(U_{i})_{i\in\mathcal{N}}$. For each guess, the algorithm runs Algorithm 1, and stores the set of EF1 contracts returned by it. Then, it outputs the contract which maximizes the principal’s utility. Let’s denote this contract with $(\tilde{S},\tilde{\alpha})$.
Now, by Proposition 4.2, Algorithm 3 finds a set $\mathcal{C}$ which includes an allocation $\hat{S}$ and a contract $\hat{\alpha}$ that induce the same utility profile as $(\bar{\alpha},S^{*})$.
Formally, we have that
$\sum_{i\in\mathcal{N}}\sum_{j\in S_{i}^{*}}\hat{U}^{P}(\bar{\alpha}_{j},i)=\sum_{i\in\mathcal{N}}\sum_{j\in\hat{S}_{i}}\hat{U}^{P}(\hat{\alpha}_{j},i)$
$\sum_{k\in S_{j}^{*}}\hat{U}_{i,k}(\bar{\alpha}_{k})=\sum_{k\in\hat{S}_{j}}\hat{U}_{i,k}(\hat{\alpha}_{k}),\quad\forall i,j\in\mathcal{N}$
$\hat{\alpha}_{j}p_{i,j}r_{j}-c_{i,j}\geq 0,\quad\forall i\in\mathcal{N},j\in\hat{S}_{i}.$
The final step is to show that $(\tilde{S},\tilde{\alpha})$ has principal’s utility at least $\rm{OPT\text{-}EF}-2\epsilon$, while by construction it is EF1 and IR. To do so, it is sufficient to show that $(\hat{S},\hat{\alpha})$ is EF1 and $2\epsilon$-optimal, since such a contract belongs to $\mathcal{C}$.
Lemma 4.7.
$(\hat{S},\hat{\alpha})$ guarantees:
$\sum_{i\in\mathcal{N}}\sum_{k\in\hat{S}_{i}}(1-\hat{\alpha}_{k})p_{i,k}r_{k}$ $\geq\rm{OPT\text{-}EF}-2\nu.$
Proof.
It holds
$\sum_{i\in\mathcal{N}}\sum_{k\in\hat{S}_{i}}(1-\hat{\alpha}_{k})p_{i,k}r_{k}$ $\geq\sum_{i\in\mathcal{N}}\sum_{k\in\hat{S}_{i}}(\hat{U}^{P}(\hat{\alpha}_{k},i)-\delta)$
$\geq\sum_{i\in\mathcal{N}}\sum_{k\in\hat{S}_{i}}\hat{U}^{P}(\hat{\alpha}_{k},i)-\nu$
$=\sum_{i\in\mathcal{N}}\sum_{k\in S^{*}_{i}}\hat{U}^{P}(\bar{\alpha}_{k},i)-\nu$
$\geq\sum_{i\in\mathcal{N}}\sum_{k\in S^{*}_{i}}(1-\bar{\alpha}_{k})p_{i,k}r_{k}-\nu$
$\geq\rm{OPT\text{-}EF}-2\nu.$
where the last inequality follows from Equation (18). ∎
Hence, we are left to show that the contract $(\hat{S},\hat{\alpha})$ is indeed EF1.
Lemma 4.8.
$(\hat{S},\hat{\alpha})$ satisfies EF1.
Proof.
Take two agents $i$ and $j$. Then
$\sum_{k\in\hat{S}_{i}}\hat{U}_{i,k}(\hat{\alpha})$ $=\sum_{k\in S^{*}_{i}}\hat{U}_{i,k}(\bar{\alpha})$ (20a)
$\geq\sum_{k\in S^{*}_{i}}\bar{\alpha}_{k}p_{i,k}r_{k}-c_{i,k}$ (20b)
$\geq\sum_{k\in S^{*}_{j}}\max\{\bar{\alpha}_{k}p_{i,k}r_{k}-c_{i,k},0\}-\nu U_{i}$ (20c)
$\geq\sum_{k\in S^{*}_{j}}\hat{U}_{i,k}(\bar{\alpha})-2\nu U_{i}$ (20d)
$=\sum_{k\in\hat{S}_{j}}\hat{U}_{i,k}(\hat{\alpha})-2\nu U_{i}$ (20e)
$\geq\sum_{k\in\hat{S}_{j}}\max\{\hat{\alpha}_{k}p_{i,k}r_{k}-c_{i,k},0\}-2\nu U_{i},$ (20f)
where (20b) is by the definition of $\hat{U}_{i,k}$, (20c) is by (17), and (20d) is by (15).
To conclude the proof, we want to show that there exists a task $k_{i,j}$ such that
$\sum_{k\in\hat{S}_{i}}\hat{\alpha}_{k}p_{i,k}r_{k}-c_{i,k}\geq\sum_{k\in\hat{S}_{j}\backslash\{k_{i,j}\}}\max\{\hat{\alpha}_{k}p_{i,k}r_{k}-c_{i,k},0\}.$
Take the task giving the largest utility $k_{i,j}\in\arg\max_{k\in\hat{S}_{j}}\hat{\alpha}_{k}p_{i,k}r_{k}-c_{i}$. Then, we have that
$\sum_{k\in\hat{S}_{i}}\hat{\alpha}_{k}p_{i,k}r_{k}-c_{i,k}$ $\geq\sum_{k\in\hat{S}_{i}}\hat{U}_{i,k}(\hat{\alpha}_{k})-\nu U_{i}$
$=(\sum_{k\in\hat{S}_{i}}\hat{U}_{i,k}(\hat{\alpha}_{k})+2\nu U_{i})-3\nu U_{i}$
$\geq(1-6\nu)(\sum_{k\in\hat{S}_{i}}\hat{U}_{i,k}(\hat{\alpha}_{k})+2\nu U_{i})$
$\geq(1-6\nu)(\sum_{k\in\hat{S}_{j}}\max\{\hat{\alpha}_{k}p_{i,k}r_{k}-c_{i,k},0\})$
$\geq\frac{1-6\nu}{1-1/m}\sum_{k\in\hat{S}_{j}\setminus k_{i,j}}\max\{\hat{\alpha}_{k}p_{i,k}r_{k}-c_{i,k},0\}$
$\geq\sum_{k\in\hat{S}_{j}\setminus k_{i,j}}\max\{\hat{\alpha}_{k}p_{i,k}r_{k}-c_{i,k},0\}$
where the first inequality is by Equation (15), the second inequality is by Equation (8) and that
$\sum_{k\in\hat{S}_{i}}\hat{U}_{i,k}(\hat{\alpha}_{k})=\sum_{k\in S_{i}^{*}}\hat{U}_{i,k}(\bar{\alpha}_{k})\geq\sum_{k\in S^{*}_{i}}\alpha_{k}^{*}p_{i,k}r_{k}-c_{i,k},$
the third inequality is by Equation (20), the fourth inequality is by the definition of $k_{i,j}$, and the last inequality is by $\nu\leq\frac{1}{6m}$. ∎
We conclude the proof noticing that the algorithm runs in polynomial time. Indeed, Algorithm 3 clearly run in polynomial time excluding the calls to Algorithm 1. Moreover, it easy to see that each call to Algorithm 1 runs in polynomial time since the possible vectors of agents/principal’s cumulative utilities have polynomial size (agent $i$ utility belongs to $\{0,\delta U_{i},\ldots,mU_{i}\}$, while the principal’s utility belongs to $\{0,\delta,\ldots,m\}$ ). ∎
5 $\nu\leftarrow\min\{\epsilon,\frac{1}{6m}$};
6 $\delta\leftarrow\nu/m$;
7 $\mathcal{C}^{*}\leftarrow\emptyset$ ;
8 for $(U_{i})_{i\in\mathcal{N}}\in\mathcal{G}^{n}$ do
9       define $\mathcal{D}_{i}$, $\mathcal{U}_{i}$, $\mathcal{H}$, $\hat{U}_{i,j}(\cdot)$, $\hat{U}^{P}(\cdot,\cdot)$ as per Equation (11), (12), (13), and (14);
10       $\mathcal{C}\leftarrow$ Run Algorithm 1 with discretization $\mathcal{D}_{i}$, $\mathcal{U}_{i}$, $\mathcal{H}$, and functions $\hat{U}_{i,j}(\cdot)$, $\hat{U}^{P}(\cdot,\cdot)$;
11       for  $(S,\alpha)\in\mathcal{C}$ do
12             if $(S,\alpha)$ satisfies EF1  then
13                   $\mathcal{C}^{*}=\mathcal{C}^{*}\cup\{{S},\alpha\}$ ;
14
15             end if
16
17       end for
18
19 end for
return $({\tilde{\alpha}},{\tilde{S}})\in\mathcal{C}^{*}$ with maximum principal utility.
2 Input: $\epsilon>0$$\{r_{j}\}_{j\in\mathcal{M}}$, $\{p_{i,j}\}_{i\in\mathcal{N},j\in\mathcal{M}}$, $\{c_{i,j}\}_{i\in\mathcal{N},j\in\mathcal{M}}$, $m,n$
Output: The contracts $(\tilde{\alpha},\tilde{S})$.
Algorithm 3 Computing EF1 Contracts
4 Input: $\epsilon>0$$\{r_{j}\}_{j\in\mathcal{M}}$, $\{p_{i,j}\}_{i\in\mathcal{N},j\in\mathcal{M}}$, $\{c_{i,j}\}_{i\in\mathcal{N},j\in\mathcal{M}}$, $m,n$
Output: The contracts $(\tilde{\alpha},\tilde{S})$.
5 Price of Fairness
[…]
If we focus on exact envy-free contracts, we show that the price of fairness is unbounded even in a case of only two agents.
Proposition 5.1.
For any constant $c>1$, there exists an instance with only two agents where the optimal revenue from envy-free contracts is less than $\frac{1}{c}$ fraction of the optimal revenue without fairness concerns, i.e., $\rm{OPT\text{-}EF}<\frac{1}{c}\cdot{\rm OPT}$.
This result can be proved with the following example.
Example 5.2.
Consider an instance parameterized by a constant $\epsilon>0$. Suppose there is one task with reward $r=1$ and two agents. Set the success probabilities and costs to $p_{1}=10\epsilon$, $c_{1}=\epsilon$ and $p_{2}=\frac{1}{2}$, $c_{2}=\frac{1}{4}$. To achieve envy-freeness, the principal must assign the task to agent $1$, and to maximize the revenue, the principal needs to offer a linear contract $\frac{c_{1}}{p_{1}}=\frac{1}{10}$ and gains a revenue of $(1-\frac{1}{10})\cdot 10\epsilon=9\epsilon$. However, without the EF constraints, the principal would assign tasks to agent $2$ and gain a revenue of $\frac{1}{4}$ by offering a linear contract $\frac{1}{2}$. The multiplicative gap in this case is $\frac{1}{36\epsilon}$, which can be arbitrarily small when $\epsilon\to 0$.
Interestingly, if we interpret the principal’s revenue as the enterprise productivity, imposing EF constraints may imply that (criticism quoted from McGinnis, (2024)) “it distorts market signals, discourages productivity…”. In the remainder of the section, we show that this gap can be improved by relaxing the EF constraints.
5.1 Relaxing EF to EF1
A natural way to reduce the gap in Proposition 5.1 is to relax the strict EF constraint. We employ a commonly adopted relaxation, EF1, and show that at least an $\frac{1}{n^{2}}$ fraction of the revenue can be extracted by EF1 contracts. This gap is small if the number of agents is not too large. For instance, in the case of two agents this gap is $1/4$, while it becomes unbounded for EF contracts.
Proposition 5.3.
For any number of agents $n$, there exists an EF1 contract with expected revenue at least $\frac{1}{n^{2}}$ fraction of the optimal revenue without fairness concerns, i.e., $\rm{OPT\text{-}EF1}\geq\frac{1}{n^{2}}\cdot{\rm OPT}$.
Proof.
Let $i^{*}$ be the agent that maximizes $\sum_{j\in\mathcal{M}}\max\{p_{i,j}r_{j}-c_{i,j},0\}$. Note that this maximum value is at least $\frac{{\rm OPT}}{n}$, where we recall that ${\rm OPT}$ is the optimal principal’s utility without fairness concerns. Set the contract for each task as $\alpha_{j}=\frac{c_{i^{*},j}}{p_{i^{*},j}r_{j}}$ if $p_{i^{*},j}r_{j}-c_{i^{*},j}\geq 0$; otherwise, set $\alpha_{j}=\min_{i\in\mathcal{P}_{j}}\frac{c_{i,j}}{p_{i,j}r_{j}}$ where $\mathcal{P}_{j}=\{i\in\mathcal{N}\mid p_{i,j}r_{j}-c_{i,j}\geq 0\}$. We recall that this set is not empty thanks to Assumption 2.3.
Now, we construct the contract using a round-robin algorithm, where agents take turns to select one of the remaining tasks. In particular, we start with agent $i^{*}$, followed by the other agents in an arbitrary order. Let $\hat{S}$ be the set of remaining tasks. In each turn, each agent $i$ chooses one of the remaining tasks in $\arg\max_{j\in\hat{S}}\alpha_{j}p_{i,j}r_{j}-c_{i,j}$ with the following additional constraints:
  • the agent selects the tasks by breaking ties to maximize the principal’s revenue;
  • the agent does not select a task if all remaining tasks are not IR.
It is straightforward to see that the modifications do not affect the conclusion that the allocation given by the round-robin algorithm is EF1. Moreover, contracts are set in order to guarantee that for any task there exists an agent that can be incentivized to exert effort. Hence, our modified round-robin algorithm can ensure that the resulting allocation is a full allocation.
Now, we provide a lower bound on the principal’s utility. To do that, it is sufficient to focus on agent $i^{*}$. Since agent $i^{*}$ has zero utility for all tasks by construction, and in each round, the tie-breaking rule assigns to $i^{*}$ the task that provides the principal’s maximum revenue (among the remaining ones), it is easy to see that the principal has revenue at least
$\frac{1}{n}\sum_{j\in\mathcal{M}}\max\{p_{i^{*},j}r_{j}-c_{i^{*},j},0\}\geq\frac{1}{n^{2}}{\rm OPT},$
implying
$\rm{OPT\text{-}EF1}\geq\frac{1}{n^{2}}\cdot{\rm OPT}.\qed$ ∎
Next, we complement the previous result by showing that the gap is at least $O(1/\sqrt{n})$. We leave as an open problem to close this gap.
Proposition 5.4.
For any number of agents $n\geq 9$, there exists an instance where the optimal revenue from EF1 contracts is at most $O(\frac{1}{\sqrt{n}})$ fraction of the optimal revenue without fairness concerns, i.e., $\rm{OPT\text{-}EF1}\leq O(\frac{1}{\sqrt{n}})\cdot{\rm OPT}$.
Proof.
Our proof takes inspiration from the construction in Bei et al., (2021). Let $m=n$ with $n\geq 9$. Also, let $r={\lfloor{\sqrt{n}}\rfloor}\geq 3$. The rewards are $1$ for all the tasks. Next, we define the success probabilities and costs for each agent.
  • For agent $i\in\{1,2,\dots,r-1\}$, let the probabilities and costs be $p_{i,(i-1)r+k}=1$, $c_{i,(i-1)r+k}=\frac{r-1}{r}$ for $k=1,2,\dots,r$, while for all other tasks $j$, let $p_{i,j}=0$, $c_{i,j}=1$.
  • For agent $i=r$, let $p_{i,j}=1$, $c_{i,j}=\frac{n-r(r-1)-1}{n-r(r-1)}$ for tasks $j\in\{r(r-1)+1,r(r-1)+2,\dots,n\}$, while for all other tasks $j$, let $p_{i,j}=0$, $c_{i,j}=1$.
  • For agent $i\in\{r+1,\dots,n\}$, let $p_{i,j}=\frac{2}{n}$, $c_{i,j}=\frac{1}{n}$ for all the task $j\in\mathcal{M}$.
Note that assigning one task to any agent $i\in\{r+1,\dots,n\}$ at most give revenue $\frac{1}{n}$ to the principal. To obtain the maximum revenue without fairness, the principal assigns to each agent $i\in\{1,2,\dots,r-1\}$ tasks $j\in\{(i-1)r+1,\dots,ir\}$ and sets the contract $\alpha_{j}=\frac{r-1}{r}>\frac{1}{n}$. The principal’s revenue from each of these contracts is $\frac{1}{r}$. Moreover, the principal assigns task $j\in\{r(r-1)+1,\dots,n\}$ to agent $r$ with contract $\alpha_{j}=\frac{n-r(r-1)-1}{n-r(r-1)}$. The revenue from each of these tasks is $\frac{1}{n-r(r-1)}>\frac{1}{n}$. This leads to a cumulative principal’s revenue of $(r-1)\cdot r\cdot\frac{1}{r}+(n-r(r-1))\cdot\frac{1}{(n-r(r-1))}=r$.
Next, we provide an upper bound for the revenue achievable with EF1 contracts. Note that by construction, agent $i\in\{1,2,\dots,r-1\}$ can only be assigned with tasks $j\in\{(i-1)r+1,\dots,ir\}$, and agent $i=r$ can only be assigned with tasks $j\in\{r(r-1)+1,\dots,n\}$.
We claim that if there exists one agent $i\in\{1,2,\dots,r\}$ assigned more than two tasks, while there is one agent $i^{\prime}\in\{r+1,\dots,n\}$ assigned no tasks, then agent $i^{\prime}$ will envy agent $i$. Note that by construction, the minimum contract to incentivize agent $i^{\prime}$ is $\frac{1}{2}$. Then, the claim holds because the minimum contract that incentivizes agent $i\in\{1,2,\dots,r-1\}$ and agent $i=r$ is greater than that for agent $i\in\{r+1,r+2,\dots,n\}$, i.e., $\frac{r-1}{r}>\frac{1}{2}$ and $\frac{n-r(r-1)-1}{n-r(r-1)}\geq\frac{r-1}{r}>\frac{1}{2}$ by $n>9$.
Recall that there are $n$ tasks. As long as there are more than two tasks assigned to one agent in $\{1,2,\dots,r\}$, each agent in $\{r+1,\dots,n\}$ must be assigned at least one task. Note that the maximum revenue from assigning task $j\in\{r(r-1)+1,\dots,n\}$ to agent $i=r$ is $\frac{1}{n-r(r-1)}<\frac{1}{r}$ since $r={\lfloor{\sqrt{n}}\rfloor}$ implies $n-r(r-1)\geq r$. Hence, the revenue of EF1 contracts in this case is at most $(n-r)\cdot\frac{1}{n}+(\frac{1}{r})\cdot(r)<2$, which is exactly the welfare extracted by assigning one task to each agent in $\{r+1,\dots,n\}$ while the remaining tasks are assigned to agents in $\{1,2,\dots,r\}$.
Moreover, if all the agents in $\{1,2,\dots,r\}$ are assigned at most one task, then the revenue of EF1 contracts is also at most the welfare $(n-r)\cdot\frac{1}{n}+(\frac{1}{r})\cdot(r)<2$. This is because assigning more tasks to agents in $\{r+1,\dots,n\}$ will lead to lower welfare, as $\frac{1}{r}\geq\frac{1}{n-r(r-1)}>\frac{1}{n}$, i.e., the maximum welfare extracted from an agent in $\{r+1,\dots,n\}$ completing one task is less than that from an agent in $\{1,2,\dots,r\}$.
In summary, the optimal revenue from EF1 contracts is at most $2$, which leads to the ratio at most $\frac{2}{r}=\frac{2}{{\lfloor{\sqrt{n}}\rfloor}}$. This concludes the proof. ∎
5.2 Achieving Fairness with Subsidies
Despite the better revenue guarantee of EF1 contracts, one potential criticism of EF1 contracts is that, while they offer some benefits, they still allows some degree of unfairness among agents. In this section, we investigate if subsidies can be used to provide better fairness guarantees while still guaranteeing a reasonable principal’s revenue. In particular, we allow the principal to compensate each agent $i$ with additional subsidies $s_{i}\geq 0$. We refer to this scheme as Envy Free with Subsidies (EFS). Under EFS, the fairness constraints require
$\sum_{k\in S_{i}}\alpha_{k}p_{i,k}r_{k}-c_{i,k}+s_{i}\geq\sum_{k\in S_{j}}\max\{\alpha_{k}p_{i,k}r_{k}-c_{i,k},0\}+s_{j},\quad\forall j\neq i$ (EFS constraints)
Note that different from EF1, EFS is exactly envy free. We revisit Example 5.2 and show that the principal revenue can be largely improved using subsidies.
Example 5.5.
Assume the same settings as Example 5.2. Let the subsidies be $s_{1}=5\epsilon$ and $s_{2}=0$. The contract is set as $\alpha=\frac{3}{5}$. By assigning the task to agent $2$, we have EFS constraint for agent $1$ as
$5\epsilon=s_{1}\geq(\alpha p_{1}r-c_{1})+s_{2}=5\epsilon$
and for agent $2$ as
$\frac{7}{20}=(\alpha p_{2}-c_{2})+s_{2}\geq s_{1}=5\epsilon$
The principal gains utility $\frac{1}{5}-5\epsilon$, much larger than $9\epsilon$ in Example 5.2.
Interestingly, while EFS contracts have more variables, the problem of computing EFS contracts is no harder than computing EF ones. We Formalize and prove this statement in Appendix B.4.
Infinite gap in EFS contracts.
The policies of compensating the weak with subsidies are indeed adopted in many countries to reduce inequality, e.g., the Temporary Assistance for Needy Families (TANF) in US. Example 5.5 seems to support that EFS is a better way to achieve fairness, as it satisfies EF and guarantees the principal’s productivity (i.e., revenue) at the same time. Unfortunately, Example 5.6 shows that this conclusion is not generally true.
Example 5.6.
Consider an instance parameterized by a constant $\epsilon>0$. Suppose there is one task with reward $r=1$ and there are $11/\epsilon+$ 1 agents. For a “special agent”, its success probability and cost are $p=\frac{1}{2}$, $c=\frac{1}{4}$. For all the other $1/\epsilon$ agents, their success probabilities and costs are $p=2\epsilon$, $c=0$. The optimal revenue in this case is to assign the task to the special agent with linear contract $\frac{1}{2}$, with expected revenue $\frac{1}{4}$.
To achieve a revenue more than $2\epsilon$, the principal must assign the task to the special agent with a contract at least $\frac{1}{2}$. However, this implies that the principal must give subsidies at least $\epsilon$ to all other agents, implying a negative principal’s revenue. Therefore, an EFS cannot obtain a revenue more than $2\epsilon$. This implies that the price of fairness is unbounded when $\epsilon\to 0$.
6 Conclusions
This paper initiates the study of a new class of fair contract design problems that capture both practical relevance and theoretical depth. Our model provides a framework for understanding how fairness constraints impact the principal’s ability to optimize profits, with direct implications for algorithmic labor platforms and other task allocation systems. By combining insights from contract theory, fair allocations, and combinatorial optimization, our work opens a new line of research at the intersection of economics and computation.
While our results establish several foundational insights, they also raise important directions for future research. In particular, we highlight two open questions that remain unresolved following our model:
  • In the case with a constant number of agents, it is unclear whether there exists a polynomial-time algorithm that achieves a constant multiplicative approximation to the optimal EF contract, under exact EF constraints or its relaxations such as $\epsilon$-EF or EF1.
  • In section 5, we provided an upper bound and a lower bound for the price of fairness in EF1 contracts. It remains an interesting question to understand which bound is indeed tight.
Beyond the specific model studied in this paper, we highlight three promising directions for future research that may be of broader interest to the community.
  • Our analysis focuses on linear contracts, which offer tractability and interpretability. A natural direction for future work is to explore nonlinear or more expressive contract classes under fairness constraints. Understanding how richer contract structures interact with fairness constraints could reveal deeper insights into the intrinsic complexity and structure of fair contract design.
  • While our objective is profit maximization, welfare maximization is an equally important goal, especially in public-sector applications such as project allocation by governments (e.g., Aharoni et al.,, 2025). In such settings, fairness remains central, but the principal may be interested in maximizing the social welfare rather than profit. Crucially, contract design in this context must still address incentive compatibility, distinguishing it from prior work on fair good allocation without strategic concerns (e.g., Bu et al.,, 2025; Aziz et al.,, 2023).
  • Our hardness results show that computing even approximately optimal contracts under envy-freeness is NP-hard. This motivates the study of alternative fairness notions that may lead to more tractable formulations. Candidates include proportionality (Suksompong,, 2016) and maximin share guarantees (Garg and Taki,, 2020), which have proven useful in other fair allocation settings and may offer a balance between fairness and computational feasibility in contract design.
Appendix A Missing Proofs
A.1 Missing Proofs for Computational Hardness
Proof of Proposition 3.5.
We prove the results by reducing from the partition problem: given a set of $\ell$ integers $n_{1},n_{2},\dots,n_{\ell}$, find a subset of integers $\mathcal{S}$ such that $\sum_{j\in\mathcal{S}}n_{j}=\sum_{j\in[\ell]}n_{j}/2$.
Construction.
There are two agents $n=2$ and $m=\ell+2$ tasks. There is a task $j^{*}$ with reward $r_{j^{*}}=1$. The success probabilities for task $j^{*}$ are $p_{1,j^{*}}=1$, $p_{2,j^{*}}=\frac{1}{10}$, and the related costs are $c_{1,j^{*}}=\frac{1}{2}$, $c_{2,j^{*}}=0$. Moreover, there is one project $\bar{j}$ with reward $r_{\bar{j}}=1$. The success probabilities relative to project $\bar{j}$ are $p_{1,\bar{j}}=\frac{1}{10},p_{2,\bar{j}}=0$, and the costs are $c_{1,\bar{j}}=c_{2,\bar{j}}=0$.
Additionally, let $T=\sum_{j\in S}n_{j}$ and $M=5T$. Then, for each $n_{j}\in\mathcal{S}$, we add a task $j$ with reward $r_{j}=\frac{n_{j}}{M}$. The success probabilities relative to this task are $p_{1,j}=1$, $p_{2,j}=\frac{1}{2}$, and the costs are $c_{1,j}=c_{2,j}=0$.
Sufficiency.
Suppose there exists a solution $\mathcal{S}_{1},\mathcal{S}_{2}$ to the partition problem, i.e., $\mathcal{S}_{1}\cup\mathcal{S}_{2}=\mathcal{S}$ and $\sum_{j\in S_{1}}n_{j}=\sum_{j\in S_{2}}n_{j}=\frac{T}{2}$. We can construct the following contracts:
  • $\alpha_{j^{*}}=\frac{1}{2}$
  • $\alpha_{j}=0$ for $j\in\mathcal{S}_{1}$
  • $\alpha_{j}=1$ for $j\in\mathcal{S}_{2}$
  • $\alpha_{\bar{j}}=1$
The tasks assigned to agent $1$ are $G_{1}=\mathcal{S}_{1}\cup\{j^{*},\bar{j}\}$, and the tasks assigned to agent $2$ are $G_{2}=\mathcal{S}_{2}$. First, by construction, we know the contracts are IR. Next, we verify that the allocations are EF. For agent $1$, the EF constraints are satisfied, i.e.,
$\sum_{j\in S_{1}}r_{j}\alpha_{j}p_{1,j}+(r_{j^{*}}\alpha_{j^{*}}p_{1,j^{*}}-c_{1,j^{*}})+r_{\bar{j}}\alpha_{\bar{j}}p_{1,\bar{j}}$ $=0+\frac{1}{10}\geq\sum_{j\in S_{2}}r_{j}\alpha_{j}p_{1,j}=\sum_{j\in S_{2}}\frac{n_{j}}{M}=\frac{\frac{1}{2}T}{5T}=\frac{1}{10}$
The EF constraints for agent $2$ are also satisfied,
$\sum_{j\in S_{2}}r_{j}\alpha_{j}p_{2,j}=\frac{1}{2}\sum_{j\in S_{2}}\frac{n_{j}}{M}=\frac{1}{20}\geq\sum_{j\in S_{1}}r_{j}\alpha_{j}p_{2,j}+(r_{j^{*}}\alpha_{j^{*}}p_{2,j^{*}}-c_{2,j^{*}})+r_{\bar{j}}\alpha_{\bar{j}}p_{2,\bar{j}}=\frac{1}{20}$
Hence, we have that the principal’s revenue is
$(1-\alpha_{j^{*}})r_{j^{*}}p_{1,j^{*}}+\sum_{j\in S_{1}}(1-\alpha_{j})r_{j}p_{1,j}=\frac{1}{2}+\sum_{j\in S_{1}}\frac{n_{j}}{M}=\frac{1}{2}+\frac{1}{10}.$
Necessity.
Next, we show that if the partition problem is not satisfiable, then the optimal revenue is smaller than $\frac{1}{2}+\frac{1}{10}$. Now, we assume that there does not exist an equal partition.
We first notice that $j^{*}$ must be assigned to agent $1$ in the optimal solution. Otherwise, the principal’s revenue is at most $\sum_{j\in\mathcal{S}}r_{j}p_{1,j}+r_{j^{*}}p_{2,j^{*}}+r_{\bar{j}}p_{1,\bar{j}}=\frac{1}{5}+\frac{1}{10}+\frac{1}{10}=\frac{2}{5}$.
Moreover, since agent $2$ has probability 00 on task $\bar{j}$, the optimal contract assigns $\bar{j}$ to agent $1$, which will not cause envy from agent $2$.
We have the following observation for the optimal solution.
Observation A.1.
In an optimal solution, the EF constraints are tight for both agents.
Proof.
Recall from the above arguments that tasks $j^{*}$ and $\bar{j}$ are assigned to agent $1$ in the optimal solution.
Suppose the optimal solution assigns $G_{1}=\mathcal{S}_{1}\cup\{j^{*},\bar{j}\}$ and $G_{2}=S_{2}$ to agent $1$ and $2$ respectively, where $\mathcal{S}_{1}\cup\mathcal{S}_{2}=\mathcal{S}$. Notice that the two agents have cost 00 for all the tasks $j\in\mathcal{S}$. This implies that in the EF constraints of a feasible solution, agent $2$ gets nonnegative utility by exerting all the tasks in $G_{1}$ (without worrying about the costs which are zero), and similarly for agent $1$. Therefore, if either one of the EF constraints is not binding in the optimal solution (e.g., the EF constraint for agent $1$), one can further decreases some contract of some task in $G_{1}$, which will increase the principal’s revenue, contradicting to the optimality. Note that such a decreasing step is always possible, as the minimum possible utility for the agent $1$ from $G_{1}$ is zero. Similar arguments also apply to agent $2$. This concludes the proof. ∎
Hence, by Observation A.1, we have the binding EF constraints as
$\sum_{j\in\mathcal{S}_{1}}r_{j}\alpha_{j}p_{1,j}+(r_{j^{*}}\alpha_{j^{*}}p_{1,j^{*}}-c_{1,j^{*}})+r_{\bar{j}}\alpha_{\bar{j}}p_{1,\bar{j}}=\sum_{j\in\mathcal{S}_{2}}r_{j}\alpha_{j}p_{1,j}$
$\sum_{j\in\mathcal{S}_{2}}r_{j}\alpha_{j}p_{2,j}=\sum_{j\in\mathcal{S}_{1}}r_{j}\alpha_{j}p_{2,j}+r_{j^{*}}\alpha_{j^{*}}p_{2,j^{*}}$
We show that in the optimal contracts, $\alpha_{j}=0$ for all $j\in\mathcal{S}_{1}$. Suppose this is not true. Then, given the optimal solution, we can construct a new solution with $\alpha_{j}^{\prime}=0$ for all $j\in\mathcal{S}_{1}$ and for all $j\in\mathcal{S}_{2}$, let $\alpha_{j}^{\prime}=\alpha_{j}-\Delta_{j}$ for some $\Delta_{j}\geq 0$ such that the EF constraints of agent $2$ is still binding,
$\sum_{j\in\mathcal{S}_{2}}r_{j}\alpha_{j}^{\prime}p_{2,j}=r_{j^{*}}\alpha_{j^{*}}p_{2,j^{*}}$ (21)
which implies that $\sum_{j\in\mathcal{S}_{2}}r_{j}\Delta_{j}p_{2,j}=\sum_{j\in\mathcal{S}_{1}}r_{j}\alpha_{j}p_{2,j}$. Due to $p_{1,j}=2p_{2,j}$ for all $j\in S$, the EF for agent $1$ is still binding,
$(r_{j^{*}}\alpha_{j^{*}}p_{1,j^{*}}-c_{1,j^{*}})+r_{\bar{j}}\alpha_{\bar{j}}p_{1,\bar{j}}=\sum_{j\in\mathcal{S}_{2}}r_{j}\alpha_{j}^{\prime}p_{1,j}$ (22)
This will lead to better principal utility, which is a contradiction. Hence, in the optimal contracts, $\alpha_{j}=0$ for all $j\in\mathcal{S}_{1}$.
To maximize the principal’s revenue, since all the utilities in $\mathcal{S}_{1}$ will be extracted by the principal ($\alpha_{j}=0$ for $j\in\mathcal{S}_{1}$), it will be optimal to maximize the principal’s revenue from tasks $j^{*}$ and $\bar{j}$, and minimize the payments in $\mathcal{S}_{2}$ to agents $2$.
Recall that $\alpha_{j^{*}}\geq\frac{1}{2}$ by IR constraint of agent $1$. By (21) and (22), we have the relation for $\alpha_{j^{*}}$ and $\alpha_{\bar{j}}$ in the optimal solution as
$\frac{4}{5}\alpha_{j^{*}}+\alpha_{\bar{j}}\frac{1}{10}=\frac{1}{2}$
Hence, to maximize the principal’s revenue from projects $j^{*}$ and $\bar{j}$, i.e.,
$(1-\alpha_{j^{*}})r_{j^{*}}p_{1,j^{*}}+(1-\alpha_{\bar{j}})r_{\bar{j}}p_{1,\bar{j}}$
we can set $\alpha_{j^{*}}=\frac{1}{2}$ and $\alpha_{\bar{j}}=1$, which indeed also minimizes the payments to agent $2$ according to (21) and $\alpha_{j^{*}}\geq\frac{1}{2}$. Hence, the optimal solution would have $\alpha_{j^{*}}=\frac{1}{2}$ and $\alpha_{\bar{j}}=1$, which implies that by (21),
$\sum_{j\in\mathcal{S}_{2}}r_{j}\alpha_{j}p_{2,j}=\frac{1}{20}$
By our previous arguments, to maximize the principal’s utility, we finally need to maximize the revenue
$\sum_{j\in\mathcal{S}_{1}}(1-\alpha_{j})r_{j}p_{1,j}+\sum_{j\in\mathcal{S}_{2}}(1-\alpha_{j})r_{j}p_{2,j}=\sum_{j\in\mathcal{S}_{1}}r_{j}p_{1,j}+\sum_{j\in\mathcal{S}_{2}}r_{j}p_{2,j}-\frac{1}{20}$
which is equivalent to maximizing $\sum_{j\in\mathcal{S}_{1}}r_{j}p_{1,j}+\sum_{j\in\mathcal{S}_{2}}r_{j}p_{2,j}$ while ensuring $\sum_{j\in\mathcal{S}_{2}}r_{j}p_{2,j}-\frac{1}{20}\geq 0$ due to that the principal revenue from agent $2$ is nonnegative.
Note that if $\sum_{j\in\mathcal{S}_{2}}r_{j}p_{2,j}=\frac{1}{20}$, it implies $\sum_{j\in\mathcal{S}_{2}}n_{j}=\frac{1}{2}T$, i.e., an equal partition. Hence, we must have
$\sum_{j\in\mathcal{S}_{2}}r_{j}p_{2,j}>\frac{1}{20}$
which due to $p_{1,j}>p_{2,j}$, implies that
$\sum_{j\in\mathcal{S}_{1}}(1-\alpha_{j})r_{j}p_{1,j}+\sum_{j\in\mathcal{S}_{2}}(1-\alpha_{j})r_{j}p_{2,j}=\sum_{j\in\mathcal{S}_{1}}r_{j}p_{1,j}+\sum_{j\in\mathcal{S}_{2}}r_{j}p_{2,j}-\frac{1}{20}<\frac{1}{10}.$
Hence, the total revenue will be
$(1-\alpha_{j^{*}})r_{j^{*}}p_{1,j^{*}}+(1-\alpha_{\bar{j}})r_{\bar{j}}p_{1,\bar{j}}+\sum_{j\in\mathcal{S}_{1}}(1-\alpha_{j})r_{j}p_{1,j}+\sum_{j\in\mathcal{S}_{2}}(1-\alpha_{j})r_{j}p_{2,j}<\frac{1}{2}+\frac{1}{10}.$
A.2 Proofs for Proposition 4.1
Proof of Proposition 4.1.
As argued, if the number of tasks $m$ is a constant, there are at most $O(n^{m})$ different allocations of tasks. To prove the result, it is sufficient to show that given an allocation $S_{1},S_{2},\dots,S_{n}$, Program (OPT-EF) is a linear program. By introducing additional variables $t_{i,k}$ for all $i\in\mathcal{N},k\in\mathcal{M}$, we have the Program (OPT-EF) become
$\max_{\{\alpha_{k}\}_{k\in\mathcal{M}}}\quad$ ${\rm Rev}(S,\alpha)$ (OPT-EF)
subject to $\alpha_{j}p_{i,j}r_{j}-c_{i,j}\geq 0,\quad\forall i\in\mathcal{N},j\in S_{i}.$
$\sum_{k\in S_{i}}\alpha_{k}p_{i,k}r_{k}-c_{i,k}\geq\sum_{k\in S_{j}}t_{i,k},\quad\forall j\neq i.$
$t_{i,k}\geq\alpha_{k}p_{i,k}r_{k}-c_{i,k},\quad\forall i\in\mathcal{N},k\in\mathcal{M}.$
$t_{i,k}\geq 0,\quad\forall i\in\mathcal{N},k\in\mathcal{M}.$
which is a linear program with polynomial number of variables and constraints. This concludes the proof. ∎
Appendix B Additional Results
B.1 Hardness for EF1
Theorem B.1.
[…]
Proof.
We prove the results by reducing from partition: Given a set $S$ of integers $\{n_{j}\}$, find a subset of elements such that their sum equals $\frac{\sum_{j\in S}n_{j}}{2}$.
Construction
There are three agents, i.e., $n=3$. Let $C=10$. Define two tasks $j^{*}$ and $\hat{j}$ with reward $r_{j^{*}}=r_{\hat{j}}=1$. The success probabilities are $p_{1,j^{*}}=p_{1,\hat{j}}=1$ and $p_{2,j^{*}}=p_{3,j^{*}}=p_{2,\hat{j}}=p_{3,\hat{j}}=\frac{1}{C}$, and the costs are $c_{1,j^{*}}=c_{1,\hat{j}}=\frac{1}{2}$ and $c_{2,j^{*}}=c_{3,j^{*}}=c_{2,\hat{j}}=c_{3,\hat{j}}=0$.
Let $M=C\sum_{j\in S}n_{j}$. For each $n_{j}$, we define a project $j$ with reward $r_{j}=\frac{n_{j}}{M}$. The success probabilities are $p_{1,j}=0$, $p_{2,j}=p_{3,j}=1$. The costs are $c_{1,j}=0$, $c_{2,j}=c_{3,j}=0$.
Sufficiency.
Suppose there is a set of indexes $S^{*}$ such that $\sum_{j\in S^{*}}n_{j}=\frac{\sum_{j\in S}n_{j}}{2}$. Assign both tasks $j^{*},\hat{j}$ to agent $1$ with $\alpha_{j^{*}}=\alpha_{\hat{j}}=1/2$. Moreover, assign all the tasks in $S^{*}$ to agent $2$, and all the other tasks in $S\setminus S^{*}$ to agent $3$. Set all the contracts $\alpha_{j}=1$ to all tasks $j\in S$.
Next, we verify that the allocation is EF1. Indeed, agent $1$ gets 00 utility from the allocated tasks, and if switching to the sets of tasks of agent $2$ or $3$, it gets utility 00.
Also, agent $2$ does not envy agent $3$, and it gets the same utility if switching to the set of tasks assigned to agent $3$. By simple calculation, we have that agent $2$ gets utility $\sum_{j\in S^{*}}\frac{n_{j}}{M}=\frac{1}{20}$. By removing the task $j^{*}$ or $\hat{j}$ from the set assigned to agent $1$, agent $2$ gets utility $\frac{1}{20}$ from the remaining task. Hence, the EF1 constraints are satisfied. A similar argument also holds for agent $3$.
Hence, the principal’s revenue (only obtained from tasks $j^{*},\hat{j}$) is $1$.
Necessity.
We show that if there exists a contract with utility strictly greater than $\frac{1}{5}+\frac{1}{2}$, then there exists a partition.
First, notice that if at most one of the tasks $j^{*},\hat{j}$ is assigned to agent $1$, the principal’s revenue is at most:
$r_{j^{*}}p_{1,j^{*}}-c_{1,j^{*}}+r_{\hat{j}}p_{2,\hat{j}}+\sum_{j\in S}r_{j}p_{2,j}=\frac{1}{2}+\frac{1}{10}+\sum_{j\in S}\frac{n_{j}}{M}=\frac{1}{5}+\frac{1}{2}.$
Hence, in order to achieve optimal utility, both tasks $j^{*}$ and $\hat{j}$ are assigned to agent $1$ and the contracts are set $\alpha_{j^{*}},\alpha_{\hat{j}}\geq 1/2$.
Since agent $2$ envies agent $1$ up to at most one task (EF1), it must be the case that $S_{2}$ (the set of tasks assigned to agent $2$) guarantees that
$\sum_{j\in S_{2}}\alpha_{j}r_{j}p_{2,j}=\sum_{j\in S_{2}}\alpha_{j}\frac{n_{j}}{M}\geq\alpha_{j^{*}}r_{j^{*}}p_{2,j^{*}}\geq\frac{1}{20}.$
which further implies
$\sum_{j\in S_{2}}\frac{n_{j}}{10\sum_{j\in S}n_{j}}=\sum_{j\in S_{2}}\frac{n_{j}}{M}\geq\frac{1}{20}$
and
$\sum_{j\in S_{2}}n_{j}\geq\frac{1}{2}\sum_{j\in S}n_{j}.$
Similar results hold for agent $3$,
$\sum_{j\in S_{3}}n_{j}\geq\frac{1}{2}\sum_{j\in S}n_{j}.$
The above together implies the existence of a partition. The gap is $\frac{7}{10}$. We conclude proof. ∎
B.2 Hardness for $\epsilon$-EF
Theorem B.2.
When there are three agents, assuming P $\neq$ NP, there does not exist a polynomial-time algorithm that computes an $\epsilon$-EF contract with revenue at least $\frac{20\epsilon+6}{10}$ fraction of ${\rm OPT\text{-}\epsilon EF}$, for any $\epsilon<\frac{1}{5}$.
Proof.
We prove the results by reducing from partition: Given a set of integers $\{n_{1},\dots,n_{\ell}\}$, find a subset of elements such that their sum equals $\frac{\sum_{j\in[\ell]}n_{j}}{2}$.
Construction
There are three agents, i.e., $n=3$. Define two tasks $j^{*}$ and $\hat{j}$ with reward $r_{j^{*}}=r_{\hat{j}}=1$. The success probabilities are $p_{1,j^{*}}=p_{1,\hat{j}}=1$ and $p_{2,j^{*}}=p_{3,j^{*}}=\frac{1}{10},p_{2,\hat{j}}=p_{3,\hat{j}}=2\epsilon$, and the costs are $c_{1,j^{*}}=c_{1,\hat{j}}=\frac{1}{2}$ and $c_{2,j^{*}}=c_{3,j^{*}}=c_{2,\hat{j}}=c_{3,\hat{j}}=0$.
Let $M=10\sum_{j\in[\ell]}n_{j}$. For each $n_{j}$, we define a project $j$ with reward $r_{j}=\frac{n_{j}}{M}$. The success probabilities are $p_{1,j}=0$, $p_{2,j}=p_{3,j}=1$. The costs are $c_{1,j}=0$, $c_{2,j}=c_{3,j}=0$.
Sufficiency.
We show that if there exists a partition, then the principal’s revenue is at least $1$.
Suppose that there is a set of indexes $\mathcal{S}$ such that $\sum_{j\in\mathcal{S}}n_{j}=\sum_{j\in[\ell]}n_{j}/2$. Consider the following contract. Assign task $j^{*}$ and $\hat{j}$ to agent $1$ with contracts as $\alpha_{j^{*}}=\alpha_{\hat{j}}=\frac{1}{2}$. Moreover, assign all the tasks in $\mathcal{S}$ to agent $2$, and all the other tasks to agent $3$. We set the contracts $\alpha_{j}=1$ to all the tasks $j\in[\ell]$.
We are left to show that the allocation satisfies the IR constraint and it guarantees revenue at least $1$. First, the allocation is $\epsilon$-EF. Indeed, by the construction above, agent $1$ receives utility 00. Due to $p_{1,j}=0$ for all $j\in[\ell]$, agent $1$ receives utility 00 switching to the tasks assigned to one of the other agents. Agent $2$ receives utility
$\sum_{j\in\mathcal{S}}r_{j}p_{2,j}-c_{2,j}=\sum_{j\in\mathcal{S}}\frac{n_{j}}{M}=\frac{1}{20}.$
Switching to the tasks assigned to agent $1$, agent $2$ gets utility
$\alpha_{j^{*}}r_{j^{*}}p_{2,j^{*}}+\alpha_{\hat{j}}r_{\hat{j}}p_{2,\hat{j}}=\frac{1}{20}+\epsilon,$
which implies it is $\epsilon$-EF. In addition, switching to the task assigned to agent $3$, agent $2$ gets utility
$\sum_{j\in[\ell]\setminus\mathcal{S}}\alpha_{j}r_{j}p_{2,j}-c_{2,j}=\frac{1}{20}$
which is due to $\mathcal{S}$ is a partition. Hence, the agent $2$ does not envy other agents. A similar argument applies to agent $3$.
We conclude noticing that the principal gets positive revenue only from task $j^{*},\hat{j}$ and in particular $1$.
Necessity.
We show that if there exists a contract with utility strictly greater than $2\epsilon+\frac{1}{10}+\frac{1}{2}$, then there exists a partition.
First, notice that if at most one of the tasks $j^{*},\hat{j}$ is assigned to agent $1$, the principal’s revenue is at most:
$r_{j^{*}}p_{1,j^{*}}-c_{1,j^{*}}+r_{\hat{j}}p_{2,\hat{j}}+\sum_{j\in[\ell]}r_{j}p_{2,j}=\frac{1}{2}+2\epsilon+\sum_{j\in[\ell]}\frac{n_{j}}{M}=2\epsilon+\frac{1}{10}+\frac{1}{2}.$
Hence, in order to achieve optimal utility, both tasks $j^{*}$ and $\hat{j}$ are assigned to agent $1$ and the contracts are set $\alpha_{j^{*}},\alpha_{\hat{j}}\geq 1/2$.
Since agent $2$ envies agent $1$ up to $\epsilon$, it must be the case that $S_{2}$ (the set of tasks assigned to agent $2$) guarantees that
$\sum_{j\in S_{2}}\alpha_{j}r_{j}p_{2,j}=\sum_{j\in S_{2}}\alpha_{j}\frac{n_{j}}{M}\geq\alpha_{j^{*}}r_{j^{*}}p_{2,j^{*}}+\alpha_{\hat{j}}r_{\hat{j}}p_{2,\hat{j}}-\epsilon\geq\frac{1}{20}.$
which further implies
$\sum_{j\in S_{2}}\frac{n_{j}}{10\sum_{j\in\{1,2,\dots,\ell\}}n_{j}}=\sum_{j\in S_{2}}\frac{n_{j}}{M}\geq\frac{1}{20}$
and
$\sum_{j\in S_{2}}n_{j}\geq\frac{1}{2}\sum_{j\in[\ell]}n_{j}.$
Similar results hold for agent $3$,
$\sum_{j\in S_{3}}n_{j}\geq\frac{1}{2}\sum_{j\in[\ell]}n_{j}.$
The above together implies the existence of a partition. The gap is $\frac{20\epsilon+6}{10}$. We conclude proof. ∎
B.3 Computing EF1 and $\epsilon$-EF Contracts with Constant Number of Tasks
Proposition B.3.
When the number of tasks is a constant, there exists a polynomial-time algorithm that computes an optimal $\epsilon$-EF and EF1 contract in ${\rm poly}(n,m)$ time.
Proof.
By proof of Proposition 4.1, the proof of $\epsilon$-EF is immediate. To prove it also holds for EF1 contracts, we note that given an allocation $S_{1},S_{2},\dots,S_{n}$, for each pair of agents $(i,j)$, the EF1 constraints can be represented as $|S_{j}|$ EF constraints by enumerating all task $g\in S_{j}$, i.e.,
$\sum_{k\in S_{i}}\alpha_{k}p_{i,k}r_{k}-c_{i,k}\geq\sum_{k\in S_{j}\backslash\{g\}}\max\{\alpha_{k}p_{i,k}r_{k}-c_{i,k},0\},\quad\forall j\neq i,\forall g\in S_{j}$
Moreover, by introducing additional variables $t_{i,k}$ for all $i\in\mathcal{N},k\in\mathcal{M}$ and rewriting the program for EF1 contracts as in the proof of Proposition 4.1, we obtain a linear program with polynomial number of variables and constraints. This concludes the proof. ∎
B.4 Reduction for Subsidies
Specifically, we prove this result by showing a reduction from EFS contracts to EF contracts.
Proposition B.4.
Given an EFS contract instance, we can reduce it to the EF contract problem in polynomial time.
Proof.
Consider the following EFS problem,
$\max\quad$ $\sum_{i\in\mathcal{N}}\Big{\{}\sum_{k\in S_{i}}(1-\alpha_{k})p_{i,k}r_{k}-s_{i}\Big{\}}$
subject to (EFS constraints), (IR constraints)
$s_{i}\geq 0,\quad\forall i\in\mathcal{N}$
Let ${\rm OPT}^{1}$ be the optimal revenue of EFS contracts. Given an EFS contracts instance, we construct the following EF contract instance: Consider a second instance with $2m+n$ tasks, where we add $m+n$ tasks (denoted as set $\mathcal{M^{+}}$) to EFS contract instance. The success probabilities and costs for any agent $i\in\mathcal{N}$ and task $k\in\mathcal{M}^{+}$ are $p_{i,k}=1$ and $c_{i,k}=0$. The reward for task $k\in\mathcal{M}^{+}$ is $r_{k}=1$.
For each agent $i$, we denote $S^{1}_{i}$ as the set of tasks assigned to agent $i$ from set $\mathcal{M}$, and as $S^{2}_{i}$ the tasks assigned to agent $i$ from $\mathcal{M}^{+}$. Hence, the program for the constructed EF contract is
$\max\quad$ $\sum_{i}[\sum_{k\in S^{1}_{i}}(1-\alpha_{k})p_{i,k}r_{k}+\sum_{k\in S^{2}_{i}}(1-\alpha_{k})p_{i,k}r_{k}]$
subject to $\sum_{k\in S^{1}_{i}}\alpha_{k}p_{i,k}r_{k}-c_{i,k}+\sum_{k\in S^{2}_{i}}\alpha_{k}p_{i,k}r_{k}\geq\sum_{k\in S^{1}_{j}}\max\{\alpha_{k}p_{i,k}r_{k}-c_{i,k},0\}+\sum_{k\in S^{2}_{j}}\alpha_{k}p_{i,k}r_{k},\quad\forall j\neq i$
(IR constraints)
Let ${\rm OPT}^{2}$ be its optimal value. Next, we show that solving the constructed EF contract instance optimally is equivalent to solving the EFS contract optimally. Specifically, we have
Claim B.5.
${\rm OPT}^{1}={\rm OPT}^{2}-m-n$.
Proof.
Given an optimal solution to the EFS program $S_{i},s_{i},\alpha_{j}$ for all $i\in\{\mathcal{N}\}$ and $j\in\{\mathcal{M}\}$, we build a solution to the EF program by setting $S^{1}_{i}=S_{i}$ and set the same contract $\alpha_{j}$ for tasks $j\in\mathcal{M}$.
First, we assign $\lceil s_{i}\rceil$ tasks from $\mathcal{M}^{+}$ to agent $i$, $S^{2}_{i}$. Additionally, we set the contract $\alpha_{j}=1$ for $\lfloor s_{i}\rfloor$ tasks in $S^{2}_{i}$. If there remains (at most) one task contract unset, set its contract as $s_{i}-\lfloor s_{i}\rfloor$. By this, we have that $\sum_{k\in S^{2}_{i}}\alpha_{k}p_{i,k}r_{k}=s_{i}$. We apply this operation to all agents $i\in\mathcal{N}$. If there still exist some tasks in $\mathcal{M}^{+}$ not assigned, assign them to an arbitrary agent with contract 00.
It is easy to see that the contract and allocation is EF, and the principal’s revenue in the constructed instance is ${\rm OPT}^{1}+m+n$.
Conversely, given an optimal solution to the constructed EF program, we can build a solution to the EFS program by taking the same allocation $S_{i}=S_{i}^{1}$ and setting the same contracts $\alpha_{j}$ for $j\in\mathcal{M}$. Additionally, set the subsidy $s_{i}=\sum_{k\in S^{2}_{i}}\alpha_{k}p_{i,k}r_{k}$ for agent $i$. It is easy to see that EFS constraints are satisfied, while the principal’s revenue decreases by $m+n$ (as we remove the added tasks). Hence, we get ${\rm OPT}^{2}-m-n$. This concludes the proof. ∎
Hence, by Claim B.5, we have that if one can solve EF contracts optimally, then one can solve EFS optimally. This prove the proposition. ∎